### v1

# Zero-shot sampling of adversarial entities in biomedical question answering

Abstract
The increasing depth of parametric domain knowledge in large language models (LLMs) is fueling their rapid deployment in real-world applications. In high-stakes and knowledge-intensive tasks, understanding model vulnerabilities is essential for quantifying the trustworthiness of model predictions and regulating their use. The recent discovery of named entities as adversarial examples in natural language processing tasks raises questions about their potential guises in other settings. Here, we propose a powerscaled distance-weighted sampling scheme in embedding space to discover diverse adversarial entities as distractors. We demonstrate its advantage over random sampling in adversarial question answering on biomedical topics. Our approach enables the exploration of different regions on the attack surface, which reveals two regimes of adversarial entities that markedly differ in their characteristics. Moreover, we show that the attacks successfully manipulate token-wise Shapley value explanations, which become deceptive in the adversarial setting. Our investigations illustrate the brittleness of domain knowledge in LLMs and reveal a shortcoming of standard evaluations for high-capacity models [^1] The code and datasets developed for the work will be made available on GitHub..
1 Introduction
LLMs such as pre-trained and finetuned generalist language models (GLMs) and their domain-adapted versions (Zhao et al., 2023; Wang et al., 2023a) incorporating transformer architectures are emerging as the technological backbone of next-generation search engines and knowledge bases (Petroni et al., 2019; Razniewski et al., 2021; Sung et al., 2021). They are increasingly used in knowledge-intensive tasks by the general public without sufficient assessment of inherent issues in trustworthiness and safety (Nat, 2023; Spitale et al., 2023; Si et al., 2023). Adversarial examples are valuable probes for model vulnerability and robustness, and are essential for performance evaluation and regulatory audits (Shayegani et al., 2023). For language models, adversarial examples are generated through guided search (Alzantot et al., 2018), sampling and probabilistic methods (Ren et al., 2019; Yang et al., 2020; Yan et al., 2022), or collected dynamically by human operators (Jia and Liang, 2017; Wallace et al., 2019b). While the search methods iteratively find random string adversarial examples (Wallace et al., 2019a), the sampling approach obviates the need for predetermined search heuristics. It can operate with only blackbox access to models and in the zero-shot setting, which is fitting for model auditing frameworks (Metaxa et al., 2021).
Figure 1: Text substitution attack on QA with adversarial distractors. (a) Illustration of the attack scheme in embedding space. A is the key to the question, B-D the original distractors, ${}^{\prime}$ and ${}^{\prime\prime}$ the adversarial distractors semantically close to and far away from the key A, respectively. (b) An example of the attack showing the correct (left) and incorrect (right) model responses (checkmarked) before and after perturbation.
The present work focuses on question answering (QA) in the biomedical domain, which has been the proving ground in recent achievements of biomedical language models (BLMs), here referring to GLMs adapted to the biomedical domain (Singhal et al., 2023a; Kung et al., 2023; Singhal et al., 2023b). GLMs have general biomedical knowledge during model development. BLMs gain further domain knowledge primarily through finetuning on a variety of text data from clinical notes (Lehman et al., 2023), electronic health records (Wornow et al., 2023), to abstracts or excerpts from scholarly publications (Tinn et al., 2023). Although QA is frequently employed for model performance evaluation (Clark and Etzioni, 2016; Gardner et al., 2019; Robinson and Wingate, 2023), currently, the performance of LLMs in biomedical QA lacks systematic evaluation and their deployment in the real world is still in its infancy (Kung et al., 2023). Specifically, model performance against adversarial examples is seldomly assessed and their efficient sampling in the biomedical domain remains challenging, leaving a gap in our understanding of model robustness in entity-rich, high-stakes domain.
Knowledge and reasoning tests in the biomedical domain assess the understanding of diverse concepts, which are represented by domain-specific named entities (NEs). In this work, we develop an in-domain adversarial probe for model knowledge on biomedical NEs through text substitution of entity mentions in the distractors of multiple-choice QA. We introduce a powerscaled distance-weighted sampling (DWS) method in embedding space to generate adversarial examples of distractors, or adversarial distractors, in the zero-shot setting. Our work reports the first use of DWS for adversarial attacks in text [^2] Throughout the text, we use DWS to refer to the class of sampling method that employs distance-dependent probability weights. The powerscaled DWS is a type of DWS, so is the other version introduced in (Wu et al., 2017).. The attack scheme enables traversing between the two regimes of adversarial entities semantically close to or far away from an anchor point (illustrated as the key A in Fig. 1a). For demonstration, we focus on attacking two of the most common NEs in the biomedical domain (Wei et al., 2019): drug and disease names, shown in an example in Fig. 1b.. In addition, we introduce the FDA-drugs and CTD-diseases datasets as the source vocabularies of the perturbation sets, along with an entity type-annotated version of existing biomedical QA datasets MedQA-USMLE (Jin et al., 2021) and MedMCQA (Pal et al., 2022). Through analysis of attack statistics, we show that our approach provides more flexibility to uniform random sampling in entity-centric attacks (Yan et al., 2022), which maps the entity-level robustness landscapes of different LLMs. We further explore the relationship between Shapley value-based model explanation (Chen et al., 2023) and adversarial attacks and uncover the defining signatures of successful attacks in this context.
2 Entity-centric adversarial attack
2.1 Adversarial distractor generation
We focus on constructing adversarial distractors in multiple-choice QA. Let each question $\mathcal{Q}$ be represented by a tuple, $\mathcal{Q}=(\mathcal{C},\mathcal{S},\mathcal{D},k)$, with the context $\mathcal{C}$, question stem $\mathcal{S}$, a distractor set $\mathcal{D}=\{d_{i}\}_{i=1}^{l}$, and a key $k$. The distractors are the wrong choices given to confuse test-takers, while the key is the correct one (CH and Saha, 2020). QA is a natural language generation task. In an autoregressive LLM, the answer corresponds to the next tokens generated following $(\mathcal{C},\mathcal{S},\mathcal{D},k)$. The adversarial distractor is a threat model in which the perturbed distractor set $\mathcal{D}^{\prime}$ takes the form
$\mathcal{D}^{\prime}=(\mathcal{D}\setminus\{d\})\cup\{d^{\prime}\},$ (1)
[…]
$\mathrm{argmax}$ $\Pr(Y=k|\mathcal{C},\mathcal{S},\mathcal{D}^{\prime},k)$
$\neq\mathrm{argmax}\Pr(Y=k|\mathcal{C},\mathcal{S},\mathcal{D},k).$ (2)
The formulation in Eq. (2) is an untargeted attack, while the targeted version only requires restricting the answer to the perturbed question to a specific distractor. We extend Eq. (2) to a more general case in Appendix A with multiple entities and additional text in the distractors. Unlike text addition attacks that introduce distracting information (Jia and Liang, 2017; Wallace et al., 2019a), the edit constraint on $\mathcal{Q}$ that is usually enforced in text adversarial attacks (Roth et al., 2021) doesn’t preclude large semantic perturbations at the entity level in the distractors. In this work, we draw perturbations from a finite perturbation set, denoted as -$\mathcal{E}_{\textrm{-}k}\coloneqq\mathcal{E}\setminus\{k\}$, with =$\mathcal{E}=\{e_{j}\}_{j=1}^{m+1}$ being the entity dataset as illustrated in Fig. 1a. In practice, $\mathcal{E}$ may be compiled from potential model use cases or curated by regulatory bodies in a model audit.
2.2 Powerscaled distance-weighted sampling
We consider sampling with distance-dependent probability weights from the embedding space of NEs. The concept of DWS was previously introduced in contrastive learning in computer vision, where the Euclidean distance between samples and an anchor point was used and the probability weights depend on the embedding dimensionality (Wu et al., 2017). We propose a more flexible version independent of the embedding dimensionality using a powerscaled distance with a tunable hyperparameter. Specifically, for a key $k$, we assign a distance-dependent probability weight $p_{j}$ to entity $e_{j}$ in -$\mathcal{E}_{\textrm{-}k}$ such that the distractor $d$ is substituted by
$d^{\prime}=e_{j}\quad\mathrm{w.p.}\quad p_{j}=\frac{h^{n}(k,e_{j})}{\sum_{j}h^{n}(k,e_{j})}.$ (3)
Here, $h(k,e_{j})$ is a distance metric between $k$ and $e_{j}$, while the exponent $n\in\mathbb{R}$ controls the shape of the probability weight distribution. When $n>0$, as $n$ increases, the entities further away from $k$ are more preferably weighted. When $n<0$, the sampling method employs inverse distance weights, the entities closer to $k$ are more preferably weighted. Besides, the formalism accommodates a deterministic regime as a limiting case [^3] The limit $n\rightarrow-\infty$ indicates replacing the distractor $d$ with $k$ at all times. Since $k$ is not in the perturbation set -$\mathcal{E}_{\textrm{-}k}$, this limit is not attained.. It corresponds to $n\rightarrow+\infty$, when $p_{j}$ is vanishingly small for all but the entity furthest from $k$ in the embedding space according to $h(k,e_{j})$,
$d^{\prime}=e_{j}\quad\mathrm{s.t.}\quad e_{j}=\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmax}}\,h(k,e_{j}).$ (4)
When $n=0$, $p_{j}$ is the same for each entity in -$\mathcal{E}_{\textrm{-}k}$, corresponding to uniform random sampling. A short proof of the limiting cases is given in Appendix B. Therefore, our approach bridges the probabilistic and deterministic approaches to sampling adversarial examples.
Figure 2: Powerscaled DWS of adversarial distractors exhibits a two-regime effect at negative and positive $n$ values (see Eq. (3)) in model accuracy (top) and diversity index (bottom) of replacement entities in successful attacks. The observed similar behaviors are compared across models and datasets in (a) Flan-T5-XXL on MedQA-USMLE, (b) MedAlpaca-7B on MedQA-USMLE, (c) Flan-T5-xl on MedMCQA, (d) Flan-UL2 on MedMCQA. Disease and drug-mention questions are separated by colors. The vertical dashed line indicates the location of random sampling.
3 Datasets
Entity datasets
To relate to real-world scenarios, we sourced the vocabulary datasets of drug and disease names from existing public databases. The drug names dataset (FDA-drugs) contains over 2.3k entities from known drugs approved by the United States Food and Drug Administration (FDA) and curated by Drug Central [^4] https://drugcentral.org/download (Ursu et al., 2017). FDA is a world authority that approves the legal distribution of drugs, therefore the dataset represents the drug-entity names encountered in daily life. The disease names dataset [^5] https://ctdbase.org/downloads/ (CTD-diseases) contains over 9.8k entities from the Comparative Toxicogenomics Database (Davis et al., 2009), which contains a comprehensive collection of uniquely documented chemical-gene-disease interactions for humans. Both databases are curated by domain experts through regular updates. The 2023 version of the datasets was used after minor data processing.
Biomedical QA datasets
We selected over 9.3k questions from the MedQA-USMLE (Jin et al., 2021) dataset and over 3.8k questions from the MedMCQA (Pal et al., 2022) dataset for benchmarking. Both datasets are entity-rich and cover a wide range of biomedical topics. The MedQA-USMLE dataset contains long-context questions that are used to assess medical students in the United States. The MedMCQA dataset contains short questions used in medical exams in India. They have been used in recent works for assessing the biomedical knowledge in LLMs (Singhal et al., 2023a; Han et al., 2023; Singhal et al., 2023b). Both datasets are publicly available and don’t contain personal information. We annotated the NEs in the two datasets according to the entity types of the Unified Medical Language System (UMLS) (Bodenreider, 2004) using scispaCy (Neumann et al., 2019). We then divided the QA datasets into drug- or disease-mention questions by the binary mention labels generated according to the UMLS entity types. (see Appendix C).
4 Attack implementation
4.1 Type-constrained entity selection
Implementing the framework in Section 2 and Appendix A for drug and disease names requires accounting for the entity types they are associated with. We used the entity type annotations according to the UMLS (Bodenreider, 2004) of the semantic groups Chemicals & Drugs and Disorders. For each question the model answers correctly, we selected the entity of the same semantic group in the distractors closest to the key as the one to attack. The choice of the closest distractor is to maintain consistency and is not uniquely associated with the effects observed here. Then, we selected the replacement entity using random sampling or powerscaled DWS from the corresponding perturbation set -$\mathcal{E}_{\textrm{-}k}$ (constructed with FDA-drugs or CTD-diseases datasets), which is equivalent to applying a type constraint to the replacement entity. For each question to attack, the probability weights in Eq. (3) were calculated for all members of -$\mathcal{E}_{\textrm{-}k}$ using the cosine distance as $h$ between the key and the replacement entities, $k$ and $e_{j}$, in a surrogate word embedding of a BERT model finetuned on biomedical text (Yuan et al., 2022),
$h(k,e_{j})=1-\frac{\text{Embed}(k)\cdot\text{Embed}(e_{j})}{\|\text{Embed}(k)\|\cdot\|\text{Embed}(e_{j})\|}.$ (5)
Using a shared embedding allows model evaluation by the same standard and simplifies the identification of transferable adversarial entities between models, as will be discussed in Section 5.
4.2 Attack execution on LLMs
We selected both generalist (GLMs) and specialist (BLMs) models for benchmarking sampling-based attacks, which require only API access to the LLM. For GLMs, we used instruction-finetuned T5 (Flan-T5) series of models (Chung et al., 2022) and UL2 model (Tay et al., 2022) (Flan-UL2), OpenAI’s GPT-3.5 (text-davinci-003) (Brown et al., 2020) and Google’s PaLM2 (text-bison-001) (Anil et al., 2023). For BLMs, we used MedAlpaca-7B (Han et al., 2023) and Writer’s Palmyra-Med-20B (Kamble and Alshikh, 2023). The selected models contain diverse encoding schemes and architectures. Some of them have appeared on performance leaderboards of related tasks.
We assessed the performance of the LLMs at the baseline level and under sampling-based attacks, all in the zero-shot setting. This resulted in our somewhat lower baseline values than the reported model accuracies that are often achieved with the help of few-shot prompting and prompt optimization to improve performance. All models were evaluated at zero temperature or in the non-sampling setting. The prompts were structured as in Appendix D. Because only up to two model queries are required for each question, the computation is much more affordable than search-based approaches, which can take thousands or more queries (Alzantot et al., 2018). For the attack based on powerscaled DWS, we tuned the hyperparameter $n$ within interval of [-50, 50] using a grid search. The interval was chosen for convenience because beyond this range, the performance drop that the attack incurred on the model doesn’t deviate much from that at or within the interval boundaries.
Drug-mention questions Disease-mention questions
 Model (size)
NoPtb
[…]
NoPtb
[…]
 Flan-T5-base (0.25B)
0.276
0.238 (13.8)
0.230 (16.7) / 0.235 (14.9)
0.260
0.226 (13.1)
0.203 (21.9) / 0.221 (15.0)
 Flan-T5-large (0.78B)
0.287
0.252 (12.2)
0.245 (14.6) / 0.243 (15.3)
0.294
0.249 (15.3)
0.230 (21.8) / 0.224 (23.8)
 Flan-T5-xl (3B)
0.318
0.285 (10.4)
0.275 (13.5) / 0.284 (10.7)
0.335
0.298 (11.0)
0.281 (16.1) / 0.293 (12.5)
 Flan-T5-xxl (11B)
0.323
0.292 (9.6)
0.279 (13.6) / 0.287 (11.1)
0.345
0.303 (12.2)
0.276 (20.0) / 0.301 (12.8)
 Flan-UL2 (20B)
0.382
0.349 (8.6)
0.335 (12.3) / 0.342 (10.5)
0.396
0.356 (10.1)
0.334 (15.7) / 0.344 (13.1)
 GPT-3.5 (175B)
0.438
0.398 (9.1)
0.389 (11.2) / 0.391 (10.7)
0.453
0.414 (8.6)
0.401 (11.5) / 0.413 (8.8)
 MedAlpaca-7B
0.460
0.431 (6.3)
0.421 (8.5) / 0.416 (9.6)
0.488
0.439 (10.0)
0.423 (13.3) / 0.436 (10.7)
 Palmyra-Med-20B
0.463
0.413 (10.8)
0.405 (12.5) / 0.414 (10.6)
0.500
0.441 (11.8)
0.425 (15.0) / 0.441 (11.8)
 PaLM2 (350B)
0.287
0.263 (8.4)
0.254 (11.5) / 0.260 (9.4)
0.328
0.302 (7.9)
0.283 (13.7) / 0.297 (11.0)
[…]
Drug-mention questions Disease-mention questions
 Model (size)
NoPtb
[…]
NoPtb
[…]
 Flan-T5-xl (3B)
0.265
0.227 (14.3)
0.213 (19.6) / 0.225 (15.1)
0.261
0.199 (23.8)
0.194 (25.7) / 0.178 (31.8)
 Flan-T5-xxl (11B)
0.291
0.254 (12.7)
0.234 (19.6) / 0.244 (16.2)
0.376
0.319 (15.2)
0.276 (26.6) / 0.286 (23.9)
 Flan-UL2 (20B)
0.416
0.358 (13.9)
0.353 (15.1) / 0.363 (12.7)
0.408
0.344 (15.7)
0.304 (25.5) / 0.336 (17.6)
 MedAlpaca-7B
0.433
0.383 (11.5)
0.381 (12.0) / 0.381 (12.0)
0.423
0.377 (10.9)
0.357 (15.6) / 0.374 (11.6)
 Palmyra-Med-20B
0.675
0.602 (10.8)
0.585 (13.3) / 0.595 (11.9)
0.661
0.574 (13.2)
0.505 (23.6) / 0.548 (17.1)
Table 2: Adversarial testing of model performance on MedMCQA dataset (same structure as Table 1)
5 Attack analysis and diagnostics
We conducted failure mode analysis using the attack statistics across models and the sampling hyperparameter and through the effects on model explainability. The following summarizes the results.
Two regimes of adversarial entities.
The embedding space picture of the substitution attack in Fig. 1a provides a geometric interpretation of powerscaled DWS to examine different types of adversarial examples by tuning $n$. We use the Gini-Simpson index (Rao, 1982), $\eta_{\text{GS}}=1-\sum_{j}q_{j}^{2}$, to quantify the diversity of adversarial examples and a low value indicates less diversity. Here, $q_{j}\in(0,1]$ is the probability of occurrence of the entity $e_{j}$ among all instances of adversarial entities obtained from sampling, and $\sum_{j}q_{j}=1$. The diversity index indicates the reusability of adversarial entities which we use as a proxy indicator for the universal adversarial attack regime (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a). By combining the accuracy and the diversity relations as $n$ varies (Fig. 2), we uncover two regimes: (i) the adversarial entities semantically close to ($n\ll 0$) the key $k$ have more diversity; (ii) the adversarial entities semantically far from ($n\gg 0$) the key $k$ are more reusable. These two regimes both lead to an increase in successful attacks. Fig. 2 shows that a larger positive $n$ in Eq. (3) samples more universal adversarial entities than in other situations, leading to a marked drop in the diversity index. We call this behavior the two-regime effect and it is observed in almost all models benchmarked on both QA datasets in varying magnitude.
Figure 3: Model robustness evaluated using R-CAP ($\Delta\text{Acc}/\text{Acc}$) on disease-mention questions from (left) MedQA-USMLE and (right) MedMCQA datasets. The horizontal order partially reflects model performance without perturbation. The colors distinguishes entity-level perturbations by different sampling methods. Perturbations by random sampling (RdPtb) are in purple. Perturbations by powerscaled DWS (DistPtb) in the $n\ll 0$ and $n\gg 0$ regimes are in blue and grey, respectively. A lower value of $\Delta\text{Acc}/\text{Acc}$ indicates a more robust model.
The two-regime effect summarizes the observation that sampling with probability weights either proportional (positive $n$) or inversely proportional (negative $n$) to distance $h$ produces adversarial entities more effectively than random sampling, as shown in Tables 1-2. The benchmarks show that the $n\ll 0$ regime tends to have a more detrimental influence on model performance than the $n\gg 0$ regime. Nevertheless, there exist prominent examples that go against the trend, such as MedAlpaca (for drug-mention questions) and Flan-T5-large in Table 1 and Flan-T5-xl (for disease-mention questions) in Table 2. At small distances, the success of the attack is related to semantic proximity (Mozes et al., 2021). Whereas at larger distances, the attack succeeds potentially due to the obscurity of the sampled entities in the question context, which may produce a profound impact on the model performance (Li et al., 2023). The reduction in the diversity index, $\eta_{\text{GS}}$, shows the dominance of a few replacement entities in successful attacks. We found the most reusable adversarial entities from ranking their occurrences in successful attacks by powerscaled DWS. For drug names, the top-ranked adversarial entities include n-acetylglucosamine (an amino sugar and anti-inflammatory drug)666n-acetylglucosamine on RxList and technetium Tc 99m exametazime (a radiopharmaceutical and contrast agent)777technetium Tc 99m exametazime on RxList. For disease names, some of the chromosome deletion syndromes and rare diseases names are more reusable than others. These entities dominate the sampling results in the $n\gg 0$ regime for all LLMs, indicating their transferability and universality. An example attack using n-acetylglucosamine as recommended medication is in Fig. 4. Adversarial entities in the $n\ll 0$ regime have improper spellings such as missing possessives or they are insufficiently detailed.
Size is not all for model robustness.
A typical metric for model robustness assessment is the change in accuracy before and after perturbation (Zhang et al., 2022). To enable comparison between datasets, we quantify the model robustness through its relative change in accuracy (Acc) upon perturbation (R-CAP), as shown partly in Fig. 3.
$\frac{\Delta\text{Acc}}{\text{Acc}}=\frac{\Delta(n_{\text{TP}}+n_{\text{TN}})}{n_{\text{TP}}+n_{\text{TN}}}$ (6)
R-CAP is equivalently the relative change in the number of correct model predictions (numbers of true positives $n_{\text{TP}}$ and true negatives $n_{\text{TN}}$) (Wang et al., 2023b). Tables 1-2 show that zero-shot adversarial sampling already incurs about 10-20% of relative change in accuracy. For the Flan-T5 series (Chung et al., 2022), R-CAP is more pronounced for smaller and less performant models, indicating the improvement of robustness as models scale up. However, the BLMs are generally more robust to entity substitution than GLMs for models of a similar size. The trend is consistent in both of the two regimes found through hyperparameter tuning of the powerscaled DWS. It is worth noting that although Palmyra-Med-20B (Kamble and Alshikh, 2023) has about three times the number of parameters than MedAlpaca-7B (Han et al., 2023), it is much more sensitive to entity perturbation. Therefore, the smaller BLM is more robust against perturbation. Another observation is that the R-CAP for models upon adversarial perturbations is asymmetrical. For most of the models, sampling adversarial entities closer to the key induces more performance drop than sampling further away.
A 73-year-old man presents to the outpatient clinic complaining of chest pain with exertion. He states that resting for a few minutes usually resolves the chest pain. Currently, he takes 81 mg of aspirin daily. He has a blood pressure of 127/85 mm Hg and heart rate of 75/min. Physical examination reveals regular heart sounds and clear lung sounds bilateral. Which medication regimen below should be added? A: Amlodipine daily. Sublingual
nitroglycerin as needed.
A: Amlodipine daily. Sublingual
nitroglycerin as needed.
B: Metoprolol and a statin daily. Sublingual
nitroglycerin as needed.   ✓
B: Metoprolol and a statin daily. Sublingual
nitroglycerin as needed.
C: Metoprolol and ranolazine
daily. Sublingual nitroglycerin as needed.
C: N-acetylglucosamine and ranolazine
daily. Sublingual nitroglycerin as needed.
D: Amlodipine and a statin daily.
Sublingual nitroglycerin as needed.
D: Amlodipine and a statin daily.
Sublingual nitroglycerin as needed.   ✓
(b) Unperturbed
(c) Adversarially perturbed
Figure 4: (Top) An entity subsitution attack using n-acetylglucosamine as the replacement entity for metoprolol creates an adversarial distractor. (Bottom) Heatmaps of token-wise Shapley values for a question before and after the adversarial attack on option C. The model prediction changes from the correct choice of B (unperturbed) to the incorrect choice of D (adversarially perturbed).
Entity substitution attacks manipulate explainability.
Post-hoc analysis is a common way to understand the rationale behind model decisions (Murdoch et al., 2019). Score-based feature importance such as that based on the Shapley values are routinely used to construct post-hoc explanation of model prediction through feature attribution (Chen et al., 2023). The stability of model explanation is a growing concern for their proper use (Alvarez-Melis and Jaakkola, 2018; Hancox-Li, 2020; Lin et al., 2023). Investigations on the susceptibility of LLM explanations to adversarial perturbation in QA are still lacking. For score-based feature importance, an explainer $\phi(\cdot)$ maps the $m$-dimensional feature $\mathbf{x}\in\mathbb{R}^{m}$ to scores $\phi(\mathbf{x})\in\mathbb{R}^{m}$. In the current context, we assess the changes in feature importance represented by Shapley values (Chen et al., 2023) before ($\phi(\mathcal{Q})$) and after ($\phi(\mathcal{Q}^{\prime})$) the adversarial attack by entity substitution $\mathcal{Q}\rightarrow\mathcal{Q}^{\prime}$. For convenience of discussion, we use Shapley profile to refer to $\phi(\mathcal{Q})$ (or $\phi(\mathcal{Q}^{\prime})$), the token-wise Shapley values calculated over $\mathcal{Q}$ (or $\mathcal{Q}^{\prime}$).
We investigated the changes in the Shapley profile in the two regimes of adversarial entities for open-source LLMs, because the calculation requires access to model weights. We found that the token with the largest Shapley value (top-1 feature) is highly correlated with the model’s prediction before and after adversarial perturbation, with an example shown in Fig. 4 produced using Flan-T5-large. This behavior is an indicator of the symbol-binding characteristic of LLMs in multiple-choice QA (Robinson and Wingate, 2023). Moreover, the Shapley profile typically exhibits pronounced changes before and after entity perturbation, especially in the text locations close to the perturbed entities (Fig. 4 and Appendix E). Together, these observations indicate that the substitution attacks effectively manipulate model explanations, because the explanation after attack aligns with the wrong prediction, as if they were correct. The behavior occurs despite unchanged context, stem, and key ($\mathcal{C}$, $\mathcal{S}$, and $k$), which should contain the most important information for understanding and answering the question (Chai and Jin, 2004).
6 Related works
Text substitution attack
Adversarial attacks in neural language models have been demonstrated from the character to the sentence level (Zhang et al., 2020). The majority of text substitution attacks employ typos (Gao et al., 2018) and synonyms (Mozes et al., 2021), which effectively impose lexical constraints. Word frequency has been used as a metric to select adversarial examples in (Mozes et al., 2021). Synonym substitution attacks have been demonstrated in sentiment analysis (Liu et al., 2022). Alzantot et al. (2018) used a genetic algorithm to search for adversarial word substitution. Model sensitivity to entity-level perturbations such as substitution and swapping create problematic situations for knowledge-intensive tasks, such as knowledge conflicts (Longpre et al., 2021), compromised performance in machine reading comprehension (Yan et al., 2022) and table interpretation tasks (Koleva et al., 2023).
Robustness in QA
The robustness of QA systems has been studied using perturbative methods (Jia and Liang, 2017) and meaning-preserving transformations (Gan and Ng, 2019; Elazar et al., 2021). Sen and Saffari (2020) introduced a set of perturbations to assess language model generalization between QA datasets. Richardson and Sabharwal (2020) proposed a distractor design method using proximity in semantic space to probe model understanding of word senses. Awadalla et al. (2022) examined the connection between in- and out-of-distribution robustness for different LLMs. The performance of LLMs is sensitive to the ordering of options (Pezeshkpour and Hruschka, 2023; Zheng et al., 2023). Overall, the brittleness of knowledge in LLMs remains a primary limitation for their applications (Razniewski et al., 2021; Elazar et al., 2021; Ji et al., 2023).
Adversarial attack characteristics
The defining characteristic of adversarial attacks is to induce false predictions in trained models. Besides, attacks may possess transferability, universality (Zou et al., 2023), and they may obstruct model explanation (Noppel and Wressnegger, 2024). Transferability refers to the effectiveness of the same example being adversarial to a number of distinct models (Papernot et al., 2016). Universal adversarial perturbations are reusable across many examples tested on the same model (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a). Model explanations may be manipulated by specifically designed perturbations to the inputs (Dombrowski et al., 2019) or biased data subsampling (Laberge et al., 2023).
7 Discussion and future work
The present study illustrates that side information introduced through entity substitution leads to noticeable degradation of model performance, by modifying only a single sampling step and without gradient computation. Moreover, type-constrained entity substitution makes sure that the replacement entity appears natural to human and is less likely to be detected using grammar check or topical filtering (Morris et al., 2020). The setting simulates realistic scenarios where the “attacks" may be initiated by unsuspecting users, such as biomedical professionals using an LLM-based clinical decision support system (Liu et al., 2023). In the examples illustrated in Fig. 4 and Appendix E, adversarial distractors can lead to disease misdiagnosis or prescription of the wrong medication, which are detrimental scenarios that may occur in algorithmic decision-making in biomedicine. Further improvement on the attack performance may be achieved by sequential sampling, an intrinsic approach to boost the sampling performance, or by changing the entity vocabulary of the perturbation set, an extrinsic approach that depends on the user and application setting. Our approach may be integrated into interactive platforms for adversarial testing in open-ended human-machine conversation (Chao et al., 2023) for online model auditing.
Future research may focus on how to make LLMs more reliable through prompt engineering (Si et al., 2023; Nori et al., 2023) or leveraging retrieval from nonparametric knowledge bases (Weikum et al., 2021; Soman et al., 2023) or text sources (Wang et al., 2023c) to improve model generalization to long-tail entities. The DWS methodology may be extended to kernel methods (Hofmann et al., 2008) or metric learning (Bellet et al., 2015), which can capture more general distance dependence. Besides vulnerability identification, the current work motivates adversarial defense strategies based on generalized distances to establish a tiered system for validating adversarial robustness in user-facing and domain-oriented applications.
Limitations
Although NEs convey parts of the key information in the question, they fall short of capturing the entirety of the textual meaning. To explore other attack surfaces within the adversarial landscape of QA, one would need to evaluate the robustness in multidimensional perturbation space (Wang et al., 2022). Besides, the meaning of Shapley values calculated here requires additional interpretation due to feature correlation. Although we use them in part to reveal the importance of specific tokens, their robustness is also worth exploring using sentence-level features (Chen et al., 2023).
Ethics statement
The scenarios considered in this work using public datasets relate to potential real-world deployment of LLMs in biomedical education and clinical decision support. Our results show that the lack of adversarial testing methods in high-stakes domains may incur serious consequences because of naturally looking adversarial examples. The design and deployment of adversarial testing should be specific to and carried out at every stage of the model development cycle, paired with tools such as adversarially robust model explainability methods for understanding model behavior.
Appendix A Adversarial distractor generation
We formulate here the general case of adversarial distractor generation when the choices contain text other than entities and one or more entities are present in a single choice. To keep the symbols consistent with Section 2, we let the question be $\mathcal{Q}=(\mathcal{C},\mathcal{S},\mathcal{D},\mathcal{K})$, with the context $\mathcal{C}$, question stem $\mathcal{S}$. We write the distractor set $\mathcal{D}=\{\mathcal{D}_{i}\}_{i=1}^{l}$, and each distractor $\mathcal{D}_{i}$ contains a set of entities $\{d_{i,j}\}_{j=1}^{l_{i}}$ and additional text, which means that
$\mathcal{D}_{i}\setminus\left(\cup_{j}\{d_{i,j}\}\right)\neq\varnothing.$ (7)
The key $\mathcal{K}$ contains the entity $k$ and other text. We use the notation $\mathcal{Q}(d_{i,j}\rightarrow d_{i,j}^{\prime})$ to denote that a distractor entity $d_{i,j}$ is substituted by $d_{i,j}^{\prime}$ in the questions $\mathcal{Q}$. If $d_{i,j}$ in the question is adversarially perturbed, the model elicits a false response, while it answers correctly to the unperturbed question $\mathcal{Q}$,
$\mathrm{argmax}$ $\Pr(Y=\mathcal{K}|\mathcal{Q}(d_{i,j}\rightarrow d_{i,j}^{\prime}))$
$\neq\mathrm{argmax}\Pr(Y=\mathcal{K}|\mathcal{Q}).$ (8)
The general case is reduced to the special case described in Section 2 when $l_{i}=1$, $\mathcal{D}_{i}=d_{i,1}\coloneqq d_{i}$, and $\mathcal{K}=k$, which recover $\mathcal{D}=\{d_{i}\}_{i=1}^{l}$, and $\mathcal{Q}=(\mathcal{C},\mathcal{S},\mathcal{D},k)$.
Appendix B Proof of the sampling regimes
Here we present a simple proof of the deterministic limit of the powerscaled distance-weighted sampling described in Eq. (3) of the main text. Although the distance function $h$ is not given a specific form, we assume it to satisfy the basic properties of a metric. Given a perturbation set -$\mathcal{E}_{\textrm{-}k}=\{e_{1},...,e_{m}\}$ with respect to a key $k$, there exists a corresponding set of distance values -$\mathcal{H}_{\textrm{-}k}$.
B.1 The limit of $n\rightarrow+\infty$
We order the elements of -$\mathcal{H}_{\textrm{-}k}$ and define the ordered set -$\mathcal{H}_{\textrm{-}k,\uparrow}=\{h_{1,\uparrow},...,h_{m,\uparrow}\}$ such that,
$h_{i,\uparrow}<h_{j,\uparrow}\quad\mathrm{iff}\quad i<j$ (9)
This indicates that $h_{m,\uparrow}=\max\,\mathcal{H}_{\textrm{-}k,\uparrow}$, and $h_{j,\uparrow}/h_{m,\uparrow}<1$ for any $j\neq m$. We compute the probability weights as $n\rightarrow+\infty$ for each entity $e_{j}\in\mathcal{E}_{\textrm{-}k}$. When $j\neq m$,
$\lim_{n\rightarrow+\infty}p_{j,\uparrow}=\lim_{n\rightarrow+\infty}\frac{h_{j,\uparrow}^{n}}{\sum_{j}h_{j,\uparrow}^{n}}$
$=\lim_{n\rightarrow+\infty}\frac{(h_{j,\uparrow}/h_{m,\uparrow})^{n}}{\sum_{j\neq m}(h_{j,\uparrow}/h_{m,\uparrow})^{n}+(h_{m,\uparrow}/h_{m,\uparrow})^{n}}$
$=\frac{0}{0+1}=0.$ (10)
Similarly, as $n\rightarrow+\infty$, the probability weight for the $m$th entity is
$\lim_{n\rightarrow+\infty}p_{m,\uparrow}=1.$ (11)
For the entities in the perturbation set -$\mathcal{E}_{\textrm{-}k}$, this indicates that only
$e_{j}=\mathrm{argmax}\,\mathcal{H}_{\textrm{-}k,\uparrow}=\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmax}}\,h(k,e_{j})$ (12)
is selected with certainty at the deterministic limit.
B.2 The limit of $n\rightarrow-\infty$
A similar approach to B.1 can be used to construct the probability weights in the limit $n\rightarrow-\infty$. As noted in the footnote of the main text, this limit is not attained since $k\notin\mathcal{E}_{\textrm{-}k}$.
B.3 The case of $n=0$
When $n=0$, it is straightforward to show that $p_{j}|_{n=0}=1/m$, regardless of $j$ using the definition in Eq. (3). This scenario therefore corresponds to uniform random sampling.
Answer the following question without explanation. [[[[Content]]]]: A 23-year-old pregnant woman at 22 weeks gestation presents with burning upon urination. She states it started 1 day ago and has been worsening despite drinking more water and taking cranberry extract. She otherwise feels well and is followed by a doctor for her pregnancy. Her temperature is 97.7°F (36.5°C), blood pressure is 122/77 mmHg, pulse is 80/min, respirations are 19/min, and oxygen saturation is 98% on room air. Physical exam is notable for an absence of costovertebral angle tenderness and a gravid uterus. [[[[Question]]]]: Which of the following are the best treatment for this patient? A: Ampicillin B: Ceftriaxone C: Doxycycline D: Nitrofurantoin [[[[Answer]]]]:
Appendix C Biomedical entity processing
The original MedQA-USMLE (Jin et al., 2021) and MedMCQA (Pal et al., 2022) datasets lack entity annotations. In the present work, we carried out a two-step annotation for both datasets.
UMLS entity types
We annotated the datasets using the type unique identifiers (TUIs) specified in the UMLS (Bodenreider, 2004) for biomedical entities. The annotation produces TUIs and the entity boundaries for entire questions and was carried out using scispaCy (Neumann et al., 2019) with the en_core_sci_md model. The annotation should be regarded as algorithmically produced, rather than human-annotated and proofread ground truth.
Binary mention labels
For entity-level perturbation, the distinction between questions with and disease entity mentions is based on the type of entities present in the distractors. We generate the binary mention labels (“Drugs?” and “Diseases?”) for each question selected from the two datasets using the UMLS semantic groups of the TUIs in (i) for the distractors. Specifically, a drug-mention question has one or more entities in the distractors of the Chemicals & Drugs semantic group. A disease-mention question has one or more entities of the Disorders semantic group in its distractors.
Appendix D Prompt formatting
We structured the prompt with bracketed tags to divide different parts (Context, Question, Answer). An example taken from the MedQA-USMLE dataset (Jin et al., 2021) is shown below. For the MedMCQA dataset (Pal et al., 2022), only the Context and Answer [^1] Using the Context tag improves model performance than the Question tag here. tags were used because the questions are short and generally contain no context. The prompt formatting improves model performance and facilitates answer parsing in the model-generated text. The increased space between paragraphs in the following is only to assist viewing and is not included in the actual prompt.
Appendix E Model evaluation and attack examples
We list some examples of successful entity substitution attacks using adversarial distractors following Fig. 4 in the main text. For each example, the text box on top (the prompt formatting removed) shows the original options on the left side and the perturbed options on the right side. The entity of interest before and after perturbation is colored in red and blue, respectively. The checkmark indicates the model prediction on unperturbed and perturbed questions. Beneath the text box are the corresponding heatmaps of token-wise Shapley values (or Shapley profile) before and after the entity substitution attack. The Shapley profiles for the MedQA-USMLE dataset (Jin et al., 2021) were generated using the model Flan-T5-large and those for the MedMCQA (Pal et al., 2022) using the model Flan-T5-xxl (Chung et al., 2022). The Shapley values were calculated using the partition explainer in shap, which uses the Owen values as an approximation (Chen et al., 2023).
Of the LLMs evaluated in this work, the proprietary models (GPT-3.5, PaLM2) were accessed through third-party APIs, while the open-source models were accessed on local cache to improve speed. The evaluations were carried out on a variety of NVIDIA GPUs depending on the model size. For smaller model such as the Flan-T5 series (Chung et al., 2022), the runtime on each type of questions (involving drug or disease mentions) in a dataset was generally between 10 and 30 minutes, while it was from one to more than four hours for larger model such Flan-UL2 (Tay et al., 2022) and Palmyra-Med-20B (Kamble and Alshikh, 2023). For proprietary models accessed through their APIs, the runtime for each type of questions is between two and four hours.
An 18-year-old man presents with a sudden loss of consciousness while playing college football. There was no history of a concussion. Echocardiography shows left ventricular hypertrophy and increased thickness of the interventricular septum. Which is the most likely pathology underlying the present condition? A: Mutation in the myosin heavy chain   ✓ A: Mutation in the myosin heavy chain B: Drug abuse B: Drug abuse C: Viral infection C: Muller barth menger syndrome   ✓ D: Autoimmunity of myocardial fibers D: Autoimmunity of myocardial fibers
(b) Unperturbed
(c) Adversarially perturbed
A 41-year-old woman comes to the physician because of a 3-month history of anxiety, difficulty falling asleep, heat intolerance, and a 6-kg (13.2-lb) weight loss. The patient’s nephew, who is studying medicine, mentioned that her symptoms might be caused by a condition that is due to somatic activating mutations of the genes for the TSH receptor. Examination shows warm, moist skin and a 2-cm, nontender, subcutaneous mass on the anterior neck. Which of the following additional findings should most raise concern for a different underlying etiology of her symptoms? A: Nonpitting edema   ✓ A: Nonpitting edema B: Atrial fibrillation B: Atrial fibrillation C: Lid lag C: Lid lag D: Fine tremor D: Nonpuerperal galactorrhea   ✓
(f) Unperturbed
(g) Adversarially perturbed
A 25-year-old man is brought to the emergency department after his girlfriend discovered him at home in a minimally responsive state. He has a history of drinking alcohol excessively and using illicit drugs. On arrival, he does not respond to commands but withdraws all extremities to pain. His pulse is 90/min, respirations are 8/min, and blood pressure is 130/90 mm Hg. Pulse oximetry while receiving bag-valve-mask ventilation shows an oxygen saturation of 95%. Examination shows cool, dry skin, with scattered track marks on his arms and legs. The pupils are pinpoint and react sluggishly to light. His serum blood glucose level is 80 mg/dL. The most appropriate next step in management is intravenous administration of which of the following? A: Naloxone A: Naloxone B: Phentolamine   ✓ B: Phentolamine C: Methadone C: Methadone   ✓ D: Naltrexone D: Tienilic acid
(j) Unperturbed
(k) Adversarially perturbed
A 70-year-old woman presents to the office for a yearly physical. She states she has recently started experiencing pain in her legs and her back. Last year, she experienced a fracture of her left arm while trying to lift groceries. The patient states that she does not consume any dairy and does not go outside often because of the pain in her legs and back. Of note, she takes carbamazepine for seizures. On exam, her vitals are within normal limits. You suspect the patient might have osteomalacia. Testing for which of the following is the next best step to confirm your suspicion? A: 25-hydroxyvitamin D   ✓ A: 25-hydroxyvitamin D B: 1,25-hydroxyvitamin D B: 1,25-hydroxyvitamin D C: Pre-vitamin D3 C: Pre-vitamin D3   ✓ D: Dietary vitamin D2 D: Stiripentol
(n) Unperturbed
(o) Adversarially perturbed
Which of the following is granted orphan drug status for the treatment of Dravet’s syndrome? A: Stiripentol A: Stiripentol B: Icatibant B: Tocilizumab   ✓ C: Pre-vitamin D3 C: Pitolisant D: Tafamidis   ✓ D: Tafamidis
(r) Unperturbed
(s) Adversarially perturbed
Which among the following may be used as a sedatives-hypnotic? A: Zolmitriptan A: Zolmitriptan   ✓ B: Zileuton B: Zileuton C: Zolpidem   ✓ C: N-acetylglucosamine D: Zalcitabine D: Zalcitabine
(v) Unperturbed
(w) Adversarially perturbed

### v2

@@
-Zero-shot sampling of adversarial entities in biomedical question answering
+Assessing biomedical knowledge robustness in large language models by query-efficient sampling attacks
 Abstract
 The increasing depth of parametric domain knowledge in large language models (LLMs) is fueling their rapid deployment in real-world applications.
-In high-stakes and knowledge-intensive tasks, understanding model vulnerabilities is essential for quantifying the trustworthiness of model predictions and regulating their use.
-The recent discovery of named entities as adversarial examples in natural language processing tasks raises questions about their potential guises in other settings.
-Here, we propose a powerscaled distance-weighted sampling scheme in embedding space to discover diverse adversarial entities as distractors.
-We demonstrate its advantage over random sampling in adversarial question answering on biomedical topics.
-Our approach enables the exploration of different regions on the attack surface, which reveals two regimes of adversarial entities that markedly differ in their characteristics.
-Moreover, we show that the attacks successfully manipulate token-wise Shapley value explanations, which become deceptive in the adversarial setting.
-Our investigations illustrate the brittleness of domain knowledge in LLMs and reveal a shortcoming of standard evaluations for high-capacity models [^1] The code and datasets developed for the work will be made available on GitHub..
+Understanding model vulnerabilities in high-stakes and knowledge-intensive tasks is essential for quantifying the trustworthiness of model predictions and regulating their use.
+The recent discovery of named entities as adversarial examples (i.e.
+adversarial entities) in natural language processing tasks raises questions about their potential impact on the knowledge robustness of pre-trained and finetuned LLMs in high-stakes and specialized domains.
+We examined the use of type-consistent entity substitution as a template for collecting adversarial entities for billion-parameter LLMs with biomedical knowledge.
+To this end, we developed an embedding-space attack based on powerscaled distance-weighted sampling to assess the robustness of their biomedical knowledge with a low query budget and controllable coverage.
+Our method has favorable query efficiency and scaling over alternative approaches based on random sampling and blackbox gradient-guided search, which we demonstrated for adversarial distractor generation in biomedical question answering.
+Subsequent failure mode analysis uncovered two regimes of adversarial entities on the attack surface with distinct characteristics and we showed that entity substitution attacks can manipulate token-wise Shapley value explanations, which become deceptive in this setting.
+Our approach complements standard evaluations for high-capacity models and the results highlight the brittleness of domain knowledge in LLMs [^1] The code and datasets developed for the work is available at https://github.com/RealPolitiX/qstab/..
 1 Introduction
-LLMs such as pre-trained and finetuned generalist language models (GLMs) and their domain-adapted versions (Zhao et al., 2023; Wang et al., 2023a) incorporating transformer architectures are emerging as the technological backbone of next-generation search engines and knowledge bases (Petroni et al., 2019; Razniewski et al., 2021; Sung et al., 2021).
-They are increasingly used in knowledge-intensive tasks by the general public without sufficient assessment of inherent issues in trustworthiness and safety (Nat, 2023; Spitale et al., 2023; Si et al., 2023).
-Adversarial examples are valuable probes for model vulnerability and robustness, and are essential for performance evaluation and regulatory audits (Shayegani et al., 2023).
-For language models, adversarial examples are generated through guided search (Alzantot et al., 2018), sampling and probabilistic methods (Ren et al., 2019; Yang et al., 2020; Yan et al., 2022), or collected dynamically by human operators (Jia and Liang, 2017; Wallace et al., 2019b).
-While the search methods iteratively find random string adversarial examples (Wallace et al., 2019a), the sampling approach obviates the need for predetermined search heuristics.
-It can operate with only blackbox access to models and in the zero-shot setting, which is fitting for model auditing frameworks (Metaxa et al., 2021).
+LLMs such as pre-trained and finetuned generalist language models (GLMs) and their domain-adapted versions (Zhao et al., 2023; Wang et al., 2023a) incorporating transformer architectures are emerging as the technological backbone of next-generation search engines and knowledge bases (Petroni et al., 2019; Sung et al., 2021).
+They are increasingly used in knowledge-intensive tasks by the general public without sufficient assessment of inherent issues in trustworthiness and safety (Anonymous, 2023; Spitale et al., 2023; Si et al., 2023).
+Adversarial examples are valuable probes for model vulnerability and robustness, and are essential for performance improvements and regulatory audits (Shayegani et al., 2023).
+Adversarial attacks are classified by query scaling behavior into low- and high-query-budget attacks (see Fig.
+1a).
+Collecting adversarial data at scale, either for adversarial training or robustness assessment, requires query-efficient attacks, first empirically studied in the vision domain (Ilyas et al., 2018; Cheng et al., 2019; Shukla et al., 2021).
+For language models, adversarial examples are generated through guided search (Alzantot et al., 2018), sampling and probabilistic methods (Ren et al., 2019; Yang et al., 2020; Yan et al., 2022), or collected dynamically by human operators (Jia & Liang, 2017; Wallace et al., 2019b).
+Most of these methods require a large amount of model evaluations (in the hundreds to thousands or beyond per instance) or whitebox access to model internals, and their evaluations tend to focus on small models (with < 1B parameters) (Maheshwary et al., 2021), which are insufficient in the era of LLMs.
+Sampling-based attacks obviate the need for predetermined search heuristics and can operate with only blackbox model access and a low query budget.
+Demonstrations of query-efficient attacks on LLMs meet an emerging demand for their continuous monitoring (Metaxa et al., 2021), which is particularly important for deploying LLMs in specialized domains.
+Question answering (QA) has been the proving ground in recent achievements of biomedical language models (BLMs), here referring to GLMs adapted to the biomedical domain (Singhal et al., 2023a; Kung et al., 2023; Singhal et al., 2023b).
+Knowledge and reasoning components in biomedical QA assess the understanding of diverse concepts, which are represented by domain-specific named entities (NEs).
+The present work focuses on the identification of adversarial entities in biomedical QA for both GLMs and BLMs.
+While GLMs have general biomedical knowledge during model development, BLMs gain further domain knowledge primarily through finetuning on a variety of text data from clinical notes (Lehman et al., 2023), electronic health records (Wornow et al., 2023), to abstracts or excerpts from scholarly publications (Tinn et al., 2023).
+Although QA is frequently employed paradigm for model performance evaluation (Gardner et al., 2019; Robinson & Wingate, 2023), currently, the performance evaluation in biomedical QA lacks robustness assessment and their deployment in the real world, albeit promising, is still in its infancy (Kung et al., 2023).
+Analyzing biomedical texts using LLMs requires them to have robust contextual knowledge and language understanding of entities composed of single or multiple words.
+Evaluating model robustness in prediction for high-stakes, specialized domains requires customized methods apart from traditional benchmarks.
+(a) (a)
+(b) (c)
+(c) (b)
 Figure 1: Text substitution attack on QA with adversarial distractors.
-(a) Illustration of the attack scheme in embedding space.
-A is the key to the question, B-D the original distractors, ${}^{\prime}$ and ${}^{\prime\prime}$ the adversarial distractors semantically close to and far away from the key A, respectively.
-(b) An example of the attack showing the correct (left) and incorrect (right) model responses (checkmarked) before and after perturbation.
-The present work focuses on question answering (QA) in the biomedical domain, which has been the proving ground in recent achievements of biomedical language models (BLMs), here referring to GLMs adapted to the biomedical domain (Singhal et al., 2023a; Kung et al., 2023; Singhal et al., 2023b).
-GLMs have general biomedical knowledge during model development.
-BLMs gain further domain knowledge primarily through finetuning on a variety of text data from clinical notes (Lehman et al., 2023), electronic health records (Wornow et al., 2023), to abstracts or excerpts from scholarly publications (Tinn et al., 2023).
-Although QA is frequently employed for model performance evaluation (Clark and Etzioni, 2016; Gardner et al., 2019; Robinson and Wingate, 2023), currently, the performance of LLMs in biomedical QA lacks systematic evaluation and their deployment in the real world is still in its infancy (Kung et al., 2023).
-Specifically, model performance against adversarial examples is seldomly assessed and their efficient sampling in the biomedical domain remains challenging, leaving a gap in our understanding of model robustness in entity-rich, high-stakes domain.
-Knowledge and reasoning tests in the biomedical domain assess the understanding of diverse concepts, which are represented by domain-specific named entities (NEs).
-In this work, we develop an in-domain adversarial probe for model knowledge on biomedical NEs through text substitution of entity mentions in the distractors of multiple-choice QA.
-We introduce a powerscaled distance-weighted sampling (DWS) method in embedding space to generate adversarial examples of distractors, or adversarial distractors, in the zero-shot setting.
+(a) Typical query scaling curves in low- and high-query-budget attack settings.
+(b) An adversarial distractor example found by type-consistent entity substitution (highlighted in blue).
+The correct (top) and incorrect (bottom) model responses (checkmarked) before and after perturbation are included.
+(c) Illustration of the attack scheme in embedding space by PDWS for the example in (b).
+$\mathcal{E}$ represents the vocabulary set.
+D is the key to the question, A-C the original distractors, C is an adversarial distractor at distance $h$ from the center at key D.
+The locations of A-C and C are relative to D in the embedding space.
+In this work, we developed an effective in-domain, query-efficient adversarial probe for the knowledge robustness of LLMs in entity-rich domains using type-consistent entity substitution (TCES).
+They operate when a large number of viable replacements ($N_{\mathcal{E}}$ > 1000) per attack exist, while only a small number of text locations are permissible.
+The setting is prevalent in specialized domains where the NEs have shared characteristics such as their type and (word or phrase) structure.
+It differs from traditional synonym substitution attacks (Liu et al., 2022), which use a much smaller number of viable replacements per attack with matching word senses but can operate on many potential text locations.
+Our focus is on generating adversarial examples of distractors (see Fig.
+1b), or adversarial distractors in biomedical QA tasks.
+For relevance to real-world scenarios (Apruzzese et al., 2023), we consider the attacks in the blackbox and hard-label setting with a limited query budget ($B\ll N_{\mathcal{E}}$).
+Our main contributions include: (i) Introduction of powerscaled distance-weighted sampling (PDWS) method in embedding space to generate adversarial entities.
 Our work reports the first use of DWS for adversarial attacks in text [^2] Throughout the text, we use DWS to refer to the class of sampling method that employs distance-dependent probability weights.
 The powerscaled DWS is a type of DWS, so is the other version introduced in (Wu et al., 2017)..
-The attack scheme enables traversing between the two regimes of adversarial entities semantically close to or far away from an anchor point (illustrated as the key A in Fig.
-1a).
-For demonstration, we focus on attacking two of the most common NEs in the biomedical domain (Wei et al., 2019): drug and disease names, shown in an example in Fig.
-1b..
-In addition, we introduce the FDA-drugs and CTD-diseases datasets as the source vocabularies of the perturbation sets, along with an entity type-annotated version of existing biomedical QA datasets MedQA-USMLE (Jin et al., 2021) and MedMCQA (Pal et al., 2022).
-Through analysis of attack statistics, we show that our approach provides more flexibility to uniform random sampling in entity-centric attacks (Yan et al., 2022), which maps the entity-level robustness landscapes of different LLMs.
-We further explore the relationship between Shapley value-based model explanation (Chen et al., 2023) and adversarial attacks and uncover the defining signatures of successful attacks in this context.
+The attack scheme has a broad coverage of the embedding space centered around an anchor point (illustrated as the key D in Fig.
+1c) and maps the entity-level robustness landscapes of different LLMs in two regimes.
+(ii) Introduction of the FDA-drugs and CTD-diseases datasets as the source vocabularies of the perturbation sets, along with an entity type-annotated version of existing biomedical QA datasets MedQA-USMLE (Jin et al., 2021) and MedMCQA (Pal et al., 2022).
+Using these data, we demonstrated the vulnerability of LLMs to two of the most common NEs in the biomedical domain (Wei et al., 2019): drug and disease names (see example in Fig.
+1b).
+(iii) We unified sampling- and gradient search-based methods in TCES as an attack template, which we used for quantitative comparison of the scaling behaviors of the attacks in the same footing.
+We found that in the blackbox setting, sampling attacks can have favorable query scaling over gradient-based attacks, when are limited by the quality of gradient estimation within a fixed query budget.
+(iv) By investigating the relationship between Shapley value-based model explanation (Chen et al., 2023) and adversarial attacks, we uncovered the defining signatures of successful attacks as a result of entity substitution.
+2 Related works
+Text substitution attack
+Adversarial attacks in neural language models have been demonstrated from the character to the sentence level (Zhang et al., 2020).
+The majority of text substitution attacks employ typos (Gao et al., 2018) and synonyms (Mozes et al., 2021), which effectively impose lexical constraints.
+Word frequency has been used as a metric to select adversarial examples in (Mozes et al., 2021).
+Synonym substitution attacks have been demonstrated in sentiment analysis (Liu et al., 2022).
+Alzantot et al.
+(2018) used a genetic algorithm to search for adversarial word substitution.
+Model sensitivity to entity-level perturbations such as substitution and swapping create problematic situations for knowledge-intensive tasks, including knowledge conflicts (Longpre et al., 2021), compromised performance in machine reading comprehension (Yan et al., 2022) and table interpretation tasks (Koleva et al., 2023).
+Robustness in QA
+The robustness of QA models has been studied using perturbation methods (Jia & Liang, 2017) and meaning-preserving transformations (Gan & Ng, 2019; Elazar et al., 2021).
+Sen & Saffari (2020) introduced a set of perturbations to assess language model generalization between QA datasets.
+Richardson & Sabharwal (2020) proposed a distractor design method using proximity in semantic space to probe model understanding of word senses.
+Awadalla et al.
+(2022) examined the connection between in- and out-of-distribution robustness for different LLMs.
+The performance of LLMs is sensitive to the ordering of options (Pezeshkpour & Hruschka, 2023; Zheng et al., 2024).
+Overall, the brittleness of knowledge in LLMs remains a primary limitation for their applications (Elazar et al., 2021; Augenstein et al., 2024).
+Adversarial attack characteristics
+The defining characteristic of adversarial attacks is to induce false predictions in trained models.
+Other characteristics include (i) their generality such as transferability, universality (Zou et al., 2023).
+Transferability refers to the effectiveness of the same example being adversarial to distinct models (Papernot et al., 2016).
+Universal adversarial perturbations are reusable across many examples tested on the same model (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a).
+(ii) The quality of adversarial examples may be judged by their perceptibility to human (Dyrmishi et al., 2023).
+(iii) Adversarial attacks can operate in the low- and high-query-budget settings (Shukla et al., 2021).
+(iv) Attacked text using specifically designed perturbations to the inputs can obstruct model explanation (Noppel & Wressnegger, 2024) and render it unuseful.
 3 Entity-centric adversarial attack
@@
 We focus on constructing adversarial distractors in multiple-choice QA.
-Let each question $\mathcal{Q}$ be represented by a tuple, $\mathcal{Q}=(\mathcal{C},\mathcal{S},\mathcal{D},k)$, with the context $\mathcal{C}$, question stem $\mathcal{S}$, a distractor set $\mathcal{D}=\{d_{i}\}_{i=1}^{l}$, and a key $k$.
-The distractors are the wrong choices given to confuse test-takers, while the key is the correct one (CH and Saha, 2020).
+Let each question $\mathcal{Q}$ be represented by a tuple, $\mathcal{Q}=(\mathcal{C},\mathcal{S},\mathcal{D},k)$, with the context $\mathcal{C}$, question stem $\mathcal{S}$ [^3] The stem in the example of Fig.
+1b is the last sentence starting with “which” on the left side.
+The context is the text before the stem., a distractor set $\mathcal{D}=\{d_{i}\}_{i=1}^{l}$, and a key $k$.
+The distractors are the wrong choices given to confuse test-takers, while the key is the correct one (CH & Saha, 2020).
 QA is a natural language generation task.
@@
 […]
-$\mathrm{argmax}$ $\Pr(Y=k|\mathcal{C},\mathcal{S},\mathcal{D}^{\prime},k)$
-$\neq\mathrm{argmax}\Pr(Y=k|\mathcal{C},\mathcal{S},\mathcal{D},k).$ (2)
+$\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=k|\mathcal{C},\mathcal{S},\mathcal{D}^{\prime},k)\neq\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=k|\mathcal{C},\mathcal{S},\mathcal{D},k).$ (2)
 The formulation in Eq.
-(2) is an untargeted attack, while the targeted version only requires restricting the answer to the perturbed question to a specific distractor.
+2 is an untargeted attack, while the targeted version only requires restricting the answer to the perturbed question to a specific distractor.
 We extend Eq.
-(2) to a more general case in Appendix A with multiple entities and additional text in the distractors.
-Unlike text addition attacks that introduce distracting information (Jia and Liang, 2017; Wallace et al., 2019a), the edit constraint on $\mathcal{Q}$ that is usually enforced in text adversarial attacks (Roth et al., 2021) doesn’t preclude large semantic perturbations at the entity level in the distractors.
+2 to a more general case in Appendix A with multiple entities and additional text in the distractors.
+Unlike text addition attacks that introduce distracting information (Jia & Liang, 2017; Wallace et al., 2019a), the edit constraint on $\mathcal{Q}$ that is usually enforced in text adversarial attacks (Roth et al., 2024) doesn’t preclude large semantic perturbations at the entity level in the distractors.
 In this work, we draw perturbations from a finite perturbation set, denoted as -$\mathcal{E}_{\textrm{-}k}\coloneqq\mathcal{E}\setminus\{k\}$, with =$\mathcal{E}=\{e_{j}\}_{j=1}^{m+1}$ being the entity dataset as illustrated in Fig.
@@
 We consider sampling with distance-dependent probability weights from the embedding space of NEs.
-The concept of DWS was previously introduced in contrastive learning in computer vision, where the Euclidean distance between samples and an anchor point was used and the probability weights depend on the embedding dimensionality (Wu et al., 2017).
+The concept of distance-weighted sampling (DWS) was previously introduced in contrastive learning in computer vision, where the Euclidean distance between samples and an anchor point was used and the probability weights depend on the embedding dimensionality (Wu et al., 2017).
 We propose a more flexible version independent of the embedding dimensionality using a powerscaled distance with a tunable hyperparameter.
-Specifically, for a key $k$, we assign a distance-dependent probability weight $p_{j}$ to entity $e_{j}$ in -$\mathcal{E}_{\textrm{-}k}$ such that the distractor $d$ is substituted by
+Specifically, for a key $k$, we assign a distance-dependent probability weight $p_{j}$ to entity $e_{j}$ such that the distractor $d$ is substituted by
 $d^{\prime}=e_{j}\quad\mathrm{w.p.}\quad p_{j}=\frac{h^{n}(k,e_{j})}{\sum_{j}h^{n}(k,e_{j})}.$ (3)
-Here, $h(k,e_{j})$ is a distance metric between $k$ and $e_{j}$, while the exponent $n\in\mathbb{R}$ controls the shape of the probability weight distribution.
+Here, $h(k,e_{j})$ is a distance metric between the anchor point $k$ and $e_{j}$, while the exponent $n\in\mathbb{R}$ controls the shape of the probability weight distribution.
 When $n>0$, as $n$ increases, the entities further away from $k$ are more preferably weighted.
-When $n<0$, the sampling method employs inverse distance weights, the entities closer to $k$ are more preferably weighted.
-Besides, the formalism accommodates a deterministic regime as a limiting case [^3] The limit $n\rightarrow-\infty$ indicates replacing the distractor $d$ with $k$ at all times.
+When $n<0$, the sampling method employs inverse distance weights, and the entities closer to $k$ are more preferably weighted.
+In the subsequent evaluations, we use $h(k,e_{j})=\texttt{CosineDist}(\mathbf{k},\mathbf{e}_{j})=1-\mathbf{k}\cdot\mathbf{e}_{j}/(\|\mathbf{k}\|\cdot\|\mathbf{e}_{j}\|)$, calculated in a text embedding of choice.
+We use here the italic letters ($k,e$) to represent text while the bolded letters ($\mathbf{k},\mathbf{e}$) are their corresponding vector representations.
+The formalism accommodates a deterministic regime as a limiting case [^4] The limit $n\rightarrow-\infty$ indicates replacing the distractor $d$ with $k$ at all times.
 Since $k$ is not in the perturbation set -$\mathcal{E}_{\textrm{-}k}$, this limit is not attained..
-It corresponds to $n\rightarrow+\infty$, when $p_{j}$ is vanishingly small for all but the entity furthest from $k$ in the embedding space according to $h(k,e_{j})$,
+It corresponds to $n\rightarrow+\infty$, when $p_{j}$ is vanishingly small for all but the entity furthest from $k$,
 $d^{\prime}=e_{j}\quad\mathrm{s.t.}\quad e_{j}=\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmax}}\,h(k,e_{j}).$ (4)
@@
 Therefore, our approach bridges the probabilistic and deterministic approaches to sampling adversarial examples.
-Figure 2: Powerscaled DWS of adversarial distractors exhibits a two-regime effect at negative and positive $n$ values (see Eq.
-(3)) in model accuracy (top) and diversity index (bottom) of replacement entities in successful attacks.
-The observed similar behaviors are compared across models and datasets in (a) Flan-T5-XXL on MedQA-USMLE, (b) MedAlpaca-7B on MedQA-USMLE, (c) Flan-T5-xl on MedMCQA, (d) Flan-UL2 on MedMCQA.
-Disease and drug-mention questions are separated by colors.
-The vertical dashed line indicates the location of random sampling.
-3 Datasets
+3.3 Sampling in zeroth-order gradient attack
+Adversarial attacks based on zeroth-order optimization (ZOO) use estimated gradients from finite difference to guide the search (Ilyas et al., 2018; Cheng et al., 2019).
+In the discrete text space, Berger et al.
+(2021) developed the DiscreteZOO attack, which contains three components: the word importance ranker, the candidate sampler, and the gradient-based optimizer.
+It operates on the word level and is validated on BERT-sized models for synonym substitution attacks.
+We discuss here the last two components, which, in each iteration, aim to find a replacement text span $e^{\prime}$ through the gradient update rule
+$\mathbf{e}^{\prime}=\mathbf{e}_{0}-\lambda\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0}),$ (5)
+where $e_{0}$ is the original text span, $\lambda$ the learning rate, and ^$\widehat{\nabla}_{\mathrm{ZO}}$ indicates the zeroth-order gradient estimator.
+We use $\mathrm{LLM}(\mathcal{Q};e_{0})$ to denote the LLM output with the unperturbed question $\mathcal{Q}$ (containing $e_{0}$) as the input, then the multi-point ($M\geq 2$) gradient estimate is constructed by querying the LLM $M$ times and computing
+^$\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0})=\mathbb{E}_{\textbf{u}}\left[\frac{\Delta\mathrm{LLM}(\mathcal{Q};e_{0})}{\Delta\mathbf{e}_{0}}\textbf{u}\right]=\frac{1}{M}\sum_{i=1,e_{i}\in\mathcal{E}_{\textrm{-}k}}^{M}\frac{\left(\mathrm{LLM}(\mathcal{Q^{\prime}};e_{i})-\mathrm{LLM}(\mathcal{Q};e_{0})\right)\cdot(\mathbf{e}_{i}-\mathbf{e}_{0})}{\|\mathbf{e}_{i}-\mathbf{e}_{0}\|^{2}}.$ (6)
+This is an example of the random directions stochastic approximation (Nesterov & Spokoiny, 2017), where $\textbf{u}=(\mathbf{e}_{j}-\mathbf{e}_{0})/\|\mathbf{e}_{j}-\mathbf{e}_{0}\|$ is a randomly oriented vector in the embedding space corresponding to the neighboring entity $e_{j}$.
+Berger et al.
+(2021) computed Eq.
+6 by random sampling of neighbors (in candidate sampler) within a similarity threshold for synonyms.
+Given a noisy gradient estimate ^$\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0})$ and the discrete nature of the search space, we can rewrite Eq.
+5 effectively as deterministic DWS in the embedding space by
+$d^{\prime}=e_{j}\quad\mathrm{s.t.}\quad e_{j}=\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmin}}\,h(e_{0},e_{j}),$ (7)
+where $h(e_{0},e_{j})=\texttt{CosineDist}(\mathbf{e}_{0}-\lambda\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0}),\mathbf{e}_{j})$, with $e_{0}$ as the anchor point, and argmin indicates the closest member within the text embedding.
+Using argmin is essential because the exact location $h(e_{0},e_{j})$ in the discrete embedding space is usually not occupied.
+While a gradient estimate of sufficient quality is essential for a successful attack, increasing $M$ will count towards the query budget.
+Therefore, in the budgeted setting, a tradeoff should exist between the efficacy of the attack and the number of queries consumed for each gradient estimation.
+4 Implementing attacks on biomedical question answering
+4.1 Dataset selection
 Entity datasets
 To relate to real-world scenarios, we sourced the vocabulary datasets of drug and disease names from existing public databases.
-The drug names dataset (FDA-drugs) contains over 2.3k entities from known drugs approved by the United States Food and Drug Administration (FDA) and curated by Drug Central [^4] https://drugcentral.org/download (Ursu et al., 2017).
+The drug names dataset (FDA-drugs) contains over 2.3k unique entities from known drugs approved by the United States Food and Drug Administration (FDA) and curated by Drug Central [^5] https://drugcentral.org/download (Ursu et al., 2017).
 FDA is a world authority that approves the legal distribution of drugs, therefore the dataset represents the drug-entity names encountered in daily life.
-The disease names dataset [^5] https://ctdbase.org/downloads/ (CTD-diseases) contains over 9.8k entities from the Comparative Toxicogenomics Database (Davis et al., 2009), which contains a comprehensive collection of uniquely documented chemical-gene-disease interactions for humans.
-Both databases are curated by domain experts through regular updates.
+The disease names dataset [^6] https://ctdbase.org/downloads/ (CTD-diseases) contains over 9.8k unique entities from the Comparative Toxicogenomics Database (Davis et al., 2009), which contains a comprehensive collection of uniquely documented chemical-gene-disease interactions for humans.
+Both databases are curated by domain experts through regular updates, and unlike traditional text corpora (Mohan & Li, 2019), the entities in these two datasets contain no redundancy.
 The 2023 version of the datasets was used after minor data processing.
@@
 We selected over 9.3k questions from the MedQA-USMLE (Jin et al., 2021) dataset and over 3.8k questions from the MedMCQA (Pal et al., 2022) dataset for benchmarking.
-Both datasets are entity-rich and cover a wide range of biomedical topics.
+Both datasets are entity-rich and cover a wide range of biomedical specialties and topics.
 The MedQA-USMLE dataset contains long-context questions that are used to assess medical students in the United States.
@@
 We annotated the NEs in the two datasets according to the entity types of the Unified Medical Language System (UMLS) (Bodenreider, 2004) using scispaCy (Neumann et al., 2019).
-We then divided the QA datasets into drug- or disease-mention questions by the binary mention labels generated according to the UMLS entity types.
-(see Appendix C).
-4 Attack implementation
-4.1 Type-constrained entity selection
-Implementing the framework in Section 2 and Appendix A for drug and disease names requires accounting for the entity types they are associated with.
-We used the entity type annotations according to the UMLS (Bodenreider, 2004) of the semantic groups Chemicals & Drugs and Disorders.
-For each question the model answers correctly, we selected the entity of the same semantic group in the distractors closest to the key as the one to attack.
-The choice of the closest distractor is to maintain consistency and is not uniquely associated with the effects observed here.
-Then, we selected the replacement entity using random sampling or powerscaled DWS from the corresponding perturbation set -$\mathcal{E}_{\textrm{-}k}$ (constructed with FDA-drugs or CTD-diseases datasets), which is equivalent to applying a type constraint to the replacement entity.
-For each question to attack, the probability weights in Eq.
-(3) were calculated for all members of -$\mathcal{E}_{\textrm{-}k}$ using the cosine distance as $h$ between the key and the replacement entities, $k$ and $e_{j}$, in a surrogate word embedding of a BERT model finetuned on biomedical text (Yuan et al., 2022),
-$h(k,e_{j})=1-\frac{\text{Embed}(k)\cdot\text{Embed}(e_{j})}{\|\text{Embed}(k)\|\cdot\|\text{Embed}(e_{j})\|}.$ (5)
-Using a shared embedding allows model evaluation by the same standard and simplifies the identification of transferable adversarial entities between models, as will be discussed in Section 5.
-4.2 Attack execution on LLMs
-We selected both generalist (GLMs) and specialist (BLMs) models for benchmarking sampling-based attacks, which require only API access to the LLM.
-For GLMs, we used instruction-finetuned T5 (Flan-T5) series of models (Chung et al., 2022) and UL2 model (Tay et al., 2022) (Flan-UL2), OpenAI’s GPT-3.5 (text-davinci-003) (Brown et al., 2020) and Google’s PaLM2 (text-bison-001) (Anil et al., 2023).
-For BLMs, we used MedAlpaca-7B (Han et al., 2023) and Writer’s Palmyra-Med-20B (Kamble and Alshikh, 2023).
-The selected models contain diverse encoding schemes and architectures.
-Some of them have appeared on performance leaderboards of related tasks.
-We assessed the performance of the LLMs at the baseline level and under sampling-based attacks, all in the zero-shot setting.
-This resulted in our somewhat lower baseline values than the reported model accuracies that are often achieved with the help of few-shot prompting and prompt optimization to improve performance.
-All models were evaluated at zero temperature or in the non-sampling setting.
-The prompts were structured as in Appendix D.
-Because only up to two model queries are required for each question, the computation is much more affordable than search-based approaches, which can take thousands or more queries (Alzantot et al., 2018).
-For the attack based on powerscaled DWS, we tuned the hyperparameter $n$ within interval of [-50, 50] using a grid search.
-The interval was chosen for convenience because beyond this range, the performance drop that the attack incurred on the model doesn’t deviate much from that at or within the interval boundaries.
-Drug-mention questions Disease-mention questions
- Model (size)
-NoPtb
+We then divided the QA datasets into drug- or disease-mention questions by the binary mention labels constructed according to the UMLS entity types (see Appendix C).
+4.2 Type-consistent entity substitution
+Implementing attacks that target entities requires accounting for type consistency and compatibility with simultaneous multi-word (or span-level) operations, because of the inseparable nature of the entity components (e.g.
+growth hormone, acquired immunodeficiency syndrome).
+Algorithm 1 Type-consistent entity substitution (TCES) template for collecting adversarial entities in distractors.
+Inputs: Question Q, LLM, entity type $\tau$, query budget $B$.
+Internals: Number of choices $N_{\text{ch}}$, text t, entity Ent, embedding Emb, correct answer k, and evaluator function $\mu$.
+Outputs: Number of queries and the replacement entity, None if unsuccessful when the query budget is exhausted.
+ TCESAttacker(Q, Model, $\tau$, $B$)
+[…]
+     for $c\leftarrow 1$ to $N_{\text{ch}}$ do
 […]
 […]
-NoPtb
 […]
 […]
- Flan-T5-base (0.25B)
-0.276
-0.238 (13.8)
-0.230 (16.7) / 0.235 (14.9)
-0.260
-0.226 (13.1)
-0.203 (21.9) / 0.221 (15.0)
- Flan-T5-large (0.78B)
-0.287
-0.252 (12.2)
-0.245 (14.6) / 0.243 (15.3)
-0.294
-0.249 (15.3)
-0.230 (21.8) / 0.224 (23.8)
- Flan-T5-xl (3B)
-0.318
-0.285 (10.4)
-0.275 (13.5) / 0.284 (10.7)
-0.335
-0.298 (11.0)
-0.281 (16.1) / 0.293 (12.5)
- Flan-T5-xxl (11B)
-0.323
-0.292 (9.6)
-0.279 (13.6) / 0.287 (11.1)
-0.345
-0.303 (12.2)
-0.276 (20.0) / 0.301 (12.8)
- Flan-UL2 (20B)
-0.382
-0.349 (8.6)
-0.335 (12.3) / 0.342 (10.5)
-0.396
-0.356 (10.1)
-0.334 (15.7) / 0.344 (13.1)
- GPT-3.5 (175B)
-0.438
-0.398 (9.1)
-0.389 (11.2) / 0.391 (10.7)
-0.453
-0.414 (8.6)
-0.401 (11.5) / 0.413 (8.8)
- MedAlpaca-7B
-0.460
-0.431 (6.3)
-0.421 (8.5) / 0.416 (9.6)
-0.488
-0.439 (10.0)
-0.423 (13.3) / 0.436 (10.7)
- Palmyra-Med-20B
-0.463
-0.413 (10.8)
-0.405 (12.5) / 0.414 (10.6)
-0.500
-0.441 (11.8)
-0.425 (15.0) / 0.441 (11.8)
- PaLM2 (350B)
-0.287
-0.263 (8.4)
-0.254 (11.5) / 0.260 (9.4)
-0.328
-0.302 (7.9)
-0.283 (13.7) / 0.297 (11.0)
-[…]
-Drug-mention questions Disease-mention questions
- Model (size)
-NoPtb
+     $i\leftarrow 0$
+     while $i<B$ do
 […]
 […]
-NoPtb
+         k $\leftarrow$ Model(Q)
 […]
 […]
- Flan-T5-xl (3B)
-0.265
-0.227 (14.3)
-0.213 (19.6) / 0.225 (15.1)
-0.261
-0.199 (23.8)
-0.194 (25.7) / 0.178 (31.8)
- Flan-T5-xxl (11B)
-0.291
-0.254 (12.7)
-0.234 (19.6) / 0.244 (16.2)
-0.376
-0.319 (15.2)
-0.276 (26.6) / 0.286 (23.9)
- Flan-UL2 (20B)
-0.416
-0.358 (13.9)
-0.353 (15.1) / 0.363 (12.7)
-0.408
-0.344 (15.7)
-0.304 (25.5) / 0.336 (17.6)
- MedAlpaca-7B
-0.433
-0.383 (11.5)
-0.381 (12.0) / 0.381 (12.0)
-0.423
-0.377 (10.9)
-0.357 (15.6) / 0.374 (11.6)
- Palmyra-Med-20B
-0.675
-0.602 (10.8)
-0.585 (13.3) / 0.595 (11.9)
-0.661
-0.574 (13.2)
-0.505 (23.6) / 0.548 (17.1)
-Table 2: Adversarial testing of model performance on MedMCQA dataset (same structure as Table 1)
-5 Attack analysis and diagnostics
-We conducted failure mode analysis using the attack statistics across models and the sampling hyperparameter and through the effects on model explainability. The following summarizes the results.
-Two regimes of adversarial entities.
+[…]
+         $i\leftarrow i+1$
+     return
+This effectively results in more substitution patterns (one to two words, two to one word, two to two words, two to three words, etc.) than the more common one-to-one synonym substitution (Mozes et al., 2021).
+We describe type-consistent entity substitution (TCES) in Algorithm 1 as the task-agnostic attack template.
+TCES is a general and contextualized attack for entities, with the goal of adversarially modifying model outputs with only one span-level substitution, implemented using token boundaries.
+The template can accommodate single or multi-query attacks and it is applicable to any task that may be formatted into QA (Gardner et al., 2019).
+[…]
+4.3 Attack execution on LLMs
+Victim LLMs
+We selected both generalist (GLMs) and specialist (BLMs) models for robustness evaluation via blackbox adversarial attacks.
+The criterion we adopted here for a model to have biomedical knowledge is that it should have a baseline performance better than random guessing (e.g.
+an accuracy of 0.25 for a specific four-choice QA task) in the relevant domain-specific task.
+For GLMs, we used instruction-finetuned T5 (Flan-T5) series of models (Chung et al., 2024) and UL2 model (Tay et al., 2022) (Flan-UL2).
+For BLMs, we used MedAlpaca-7B (Han et al., 2023), Llama2-Medtuned-13B (LM2-Medtuned-13B) (Rohanian et al., 2024) and Writer’s Palmyra-Med-20B (Kamble & Alshikh, 2023).
+The selected models contain diverse encoding schemes and architectures and their sizes are $\sim$ 1-20B.
+Attack settings
+To single out the effect of the attacks, all models were evaluated at zero temperature or in the non-sampling setting and model inference was conducted in the zero-shot setting with only basic prompt instructions (see the prompt structure in Appendix D).
+Model inference of Palmyra-Med-20B (Kamble & Alshikh, 2023) used 4-bit quantization to improve speed (Dettmers & Zettlemoyer, 2023).
+The evaluation settings resulted in our somewhat lower baseline accuracies than the reported ones which are often achieved with the help of few-shot prompting and prompt optimization.
+The DiscreteZOO attack (Berger et al., 2021) was originally implemented in the textattack framework (Morris et al., 2020b), which we updated to be compatible with span-level perturbation using its boundaries, multi-GPU split-model inference, and decoder-based billion-parameter LLMs by modifying the attack recipe.
+To demonstrate the generality of the attack settings, we focus on the following three aspects in the benchmarks.
+  • Query budget: We used a fixed query budget for three types of model evaluations: (i) Single-query (budget = 1) sampling-based attacks were used as the reference because the DiscreteZOO attack requires a minimum of 3 model queries; (ii) Multi-query attacks used a budget of 8 for reasonable computational cost across all models and attack settings for both sampling- and search-based attacks; (iii) The query scaling trends of specific LLMs and different attack settings were investigated with fixed budgets under 100 per input instance.
+  • Text embedding: Both PDWS and the DiscreteZOO attacks require sampling from a text embedding.
+The counter-fitted GloVe embedding (Mrkšić et al., 2016) was a default for DiscreteZOO, but GloVe is poor for representing specialized vocabulary in the biomedical domain (Wang et al., 2018).
+Therefore, here we chose two recent embedding models that offer an improved latent representation of text relations for the benchmarks: (i) CODER (Yuan et al., 2022), a BERT-structured embedding model contrastively pretrained to distinguish UMLS knowledge graph concepts; (ii) GTE-base (Li et al., 2023b) from constrastive learning of general web-scale text data.
+  • Hyperparameters: For single-query PDWS attacks, we tuned the primary hyperparameter $n$ within the interval of [-50, 50] using grid search with a step of 5 or 10 because the local maxima of ASR appear on the positive and negative sides.
+For multi-query PDWS attacks, hyperparameter $n$ was briefly tre-uned around its optimal value in the single-query attack.
+Most tuned hyperparameters fall within [-30, -5] and [5, 30].
+Operating the DiscreteZOO attack in a compatible setting for entities requires two key modifications: (i) Use a random word ranking, which allows the attack algorithm to operate in the low-query-budget setting.
+(ii) Disable the similarity threshold for candidate point selection in gradient estimation because only entity type consistency is required.
+The ASR was greatly improved after this step.
+Benchmarks of the DiscreteZOO attacks were carried out with these settings.
+5 Finding and assessing adversarial entities
+We conducted failure mode analysis using the attack statistics gathered across models and hyperparameter settings.
+The following analyses compare methods to identify adversarial entities through query-efficient attacks and uncover their characteristics, including the distribution in the embedding space and their effects on model explainability.
+Figure 2: Powerscaled DWS of adversarial distractors exhibits a two-regime effect at negative and positive $n$ values (see Eq.
+3) in model accuracy (top) and diversity index (bottom) of replacement entities in successful attacks.
+The observed similar behaviors are compared across models and datasets in (a) Flan-T5-xxl on MedQA-USMLE, (b) Palmyra-Med-20B on MedQA-USMLE, (c) Flan-T5-xl on MedMCQA, (d) Flan-UL2 on MedMCQA.
+Disease and drug-mention questions are separated by colors.
+The vertical dashed line indicates the location of random sampling.
+{NiceTabular}
+@|l @ P1cm @ |P1.2cm @ P1.2cm @ P2.4cm @ |P1.2cm @ P1.2cm @ P2.4cm |@ MedQA-USMLE MedMCQA
+Model (size) Entity
+mention Baseline
+(Acc) RandS@@@@1
+(ASR %) PDWS@@@@1 (-/+)
++CODER (ASR %) Baseline
+(Acc) RandS@@@@1
+(ASR %) PDWS@@@@1 (-/+)
++CODER (ASR %)
+Flan-T5-xl (3B) Drugs 0.318 10.4 13.5 / 10.7 0.265 14.3 19.6 / 15.1
+Diseases 0.335 11.0 16.1 / 12.5 0.261 23.8 25.7 / 31.8
+Flan-T5-xxl (11B) Drugs 0.323 9.6 13.6 / 11.2 0.291 12.7 19.6 / 16.2
+Diseases 0.345 12.2 20.0 / 12.8 0.376 15.2 26.6 / 23.9
+Flan-UL2 (20B) Drugs 0.382 8.6 12.3 / 10.5 0.416 13.9 15.1 / 12.7
+Diseases 0.396 10.1 15.7 / 13.1 0.408 15.7 25.5 / 17.7
+MedAlpaca-7B Drugs 0.474 7.4 9.3 / 10.3 0.433 11.6 13.2 / 12.7
+Diseases 0.488 9.9 13.3 / 10.7 0.436 13.5 18.1 / 14.2
+LM2-Medtuned-13B Drugs 0.371 16.4 18.1 / 17.8 0.568 15.3 17.3 / 18.1
+Diseases 0.398 19.6 23.1 / 19.6 0.515 19.0 27.6 / 21.0
+Palmyra-Med-20B Drugs 0.382 12.6 13.3 / 13.5 0.614 10.2 13.8 / 12.7
+Diseases 0.441 16.9 21.5 / 18.7 0.624 15.4 20.1 / 11.9
+Table 1: Single-query (budget = 1) attack success rates (ASRs) on subsets of MedQA-USMLE and MedMCQA datasets by perturbing the drug or disease entity mentions.
+Zero-shot baseline (Baseline) accuracy is included for reference next to the ASR of sampling attacks using perturbation by random sampling (RandS), powerscaled distance-weighted sampling (PDWS).
+The “-” and “+” signs indicate the two regimes ($n<0$ and $n>0$).
+For each model, the highest ASR for a type of entity-mention question in each dataset is bolded.
+5.1 Two regimes of adversarial entities
+Diversity quantification of adversarial entities
 The embedding space picture of the substitution attack in Fig.
@@
 The diversity index indicates the reusability of adversarial entities which we use as a proxy indicator for the universal adversarial attack regime (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a).
-By combining the accuracy and the diversity relations as $n$ varies (Fig.
-2), we uncover two regimes: (i) the adversarial entities semantically close to ($n\ll 0$) the key $k$ have more diversity; (ii) the adversarial entities semantically far from ($n\gg 0$) the key $k$ are more reusable.
-These two regimes both lead to an increase in successful attacks.
+The dependence of the ASR and diversity on the hyperparameter $n$ in Fig.
+2 indicates two regimes: (i) the adversarial entities semantically close to ($n<0$) the key $k$ have a greater diversity; (ii) the adversarial entities semantically far from ($n>0$) the key $k$ are more reusable.
+TCES in these two regimes both can lead to an increase in successful attacks from random sampling (equivalent to $n=0$), as shown in Tables 5-5.1.
 Fig.
 2 shows that a larger positive $n$ in Eq.
-(3) samples more universal adversarial entities than in other situations, leading to a marked drop in the diversity index.
-We call this behavior the two-regime effect and it is observed in almost all models benchmarked on both QA datasets in varying magnitude.
-Figure 3: Model robustness evaluated using R-CAP ($\Delta\text{Acc}/\text{Acc}$) on disease-mention questions from (left) MedQA-USMLE and (right) MedMCQA datasets.
-The horizontal order partially reflects model performance without perturbation.
-The colors distinguishes entity-level perturbations by different sampling methods.
-Perturbations by random sampling (RdPtb) are in purple.
-Perturbations by powerscaled DWS (DistPtb) in the $n\ll 0$ and $n\gg 0$ regimes are in blue and grey, respectively.
-A lower value of $\Delta\text{Acc}/\text{Acc}$ indicates a more robust model.
-The two-regime effect summarizes the observation that sampling with probability weights either proportional (positive $n$) or inversely proportional (negative $n$) to distance $h$ produces adversarial entities more effectively than random sampling, as shown in Tables 1-2.
-The benchmarks show that the $n\ll 0$ regime tends to have a more detrimental influence on model performance than the $n\gg 0$ regime.
-Nevertheless, there exist prominent examples that go against the trend, such as MedAlpaca (for drug-mention questions) and Flan-T5-large in Table 1 and Flan-T5-xl (for disease-mention questions) in Table 2.
-At small distances, the success of the attack is related to semantic proximity (Mozes et al., 2021).
-Whereas at larger distances, the attack succeeds potentially due to the obscurity of the sampled entities in the question context, which may produce a profound impact on the model performance (Li et al., 2023).
-The reduction in the diversity index, $\eta_{\text{GS}}$, shows the dominance of a few replacement entities in successful attacks.
-We found the most reusable adversarial entities from ranking their occurrences in successful attacks by powerscaled DWS.
-For drug names, the top-ranked adversarial entities include n-acetylglucosamine (an amino sugar and anti-inflammatory drug)666n-acetylglucosamine on RxList and technetium Tc 99m exametazime (a radiopharmaceutical and contrast agent)777technetium Tc 99m exametazime on RxList.
+3 samples more universal adversarial entities than in other situations, leading to a marked drop in the diversity index.
+Moreover, there exist local maxima of ASR in both regimes with finite $n$, which were discovered by hyperparameter tuning.
+We call these behaviors the two-regime effect and they were observed in almost all models evaluated here in varying magnitudes for both QA datasets, thanks to the controllable spatial coverage of sampling methods incorporating distance information.
+In both single-query and multi-query attacks, the benchmarks show that the $n<0$ regime of PDWS attacks tends to have a greater influence on model performance than the $n>0$ regime, although prominent counterexamples do exist (MedAlpaca-7B and Llama2-Medtuned-13B (for drug-mention questions), and Flan-T5-xl (for disease-mention questions) in Table 5).
+Figure 3: Model robustness from single-query adversarial attacks using results in Table 5.
+The models (GLMs and BLMs) were evaluated on drug-mention questions from MedQA-USMLE (left) and disease-mention questions from MedMCQA (right) datasets.
+The GLMs and BLMs are ordered horizontally by their sizes.
+The bar colors distinguish between different attack methods.
+Perturbations by random sampling (RS) are in purple.
+Perturbations by powerscaled DWS (PDWS) in the $n<0$ and $n>0$ regimes are in blue and grey, respectively.
+Adversarial entity characteristics
+The two-regime effect indicates that the success of the attack may come from two sources.
+At small distances, semantic proximity (Mozes et al., 2021) reduces the distinguishability between the key and the distractor, therefore increasing the chance of model failure.
+Whereas at larger distances, the attack succeeds potentially due to the obscurity of the distractor entities in the question context, which can also profoundly impact the model performance (Li et al., 2023a).
+This explanation may be further supported by examining the diversity index, $\eta_{\text{GS}}$, which shows the dominance of a few replacement entities in successful attacks at large distances.
+We found these highly reusable adversarial entities by ranking their occurrences in successful attacks by powerscaled DWS.
+For drug names, the top-ranked adversarial entities include n-acetylglucosamine (an amino sugar and anti-inflammatory drug)777n-acetylglucosamine on RxList and technetium Tc 99m exametazime (a radiopharmaceutical and contrast agent)888technetium Tc 99m exametazime on RxList.
 For disease names, some of the chromosome deletion syndromes and rare diseases names are more reusable than others.
-These entities dominate the sampling results in the $n\gg 0$ regime for all LLMs, indicating their transferability and universality.
-An example attack using n-acetylglucosamine as recommended medication is in Fig.
-4.
-Adversarial entities in the $n\ll 0$ regime have improper spellings such as missing possessives or they are insufficiently detailed.
-Size is not all for model robustness.
-A typical metric for model robustness assessment is the change in accuracy before and after perturbation (Zhang et al., 2022).
-To enable comparison between datasets, we quantify the model robustness through its relative change in accuracy (Acc) upon perturbation (R-CAP), as shown partly in Fig.
+These entities dominate the sampling results in the $n>0$ regime for all LLMs, indicating their transferability and universality.
+{NiceTabular}
+@|l @ P1cm @ |P1cm @ P1.8cm @ P1.3cm @ P1.3cm @ |P1cm @ P1.8cm @ P1.3cm @ P1.3cm |@ MedQA-USMLE (ASR %) MedMCQA (ASR %)
+Model (size) Entity
+mention RandS
+@@@@8 PDWS@@@@8 (-/+)
++CODER DZOO@@@@8
++CODER DZOO@@@@8
++GTE-base RandS
+@@@@8 PDWS@@@@8 (-/+)
++CODER DZOO@@@@8
++CODER DZOO@@@@8
++GTE-base
+Flan-T5-xl (3B) Drugs 22.0 29.3 / 23.3 21.7 19.5 32.5 48.3 / 36.2 34.3 28.2
+Diseases 27.2 38.8 / 24.8 20.9 23.2 47.1 61.7 / 49.0 45.8 39.3
+Flan-UL2 (20B) Drugs 22.5 27.5 / 22.3 20.4 19.2 26.4 35.6 / 22.4 30.7 26.4
+Diseases 25.8 34.9 / 26.8 25.7 24.0 39.2 49.5 / 31.6 41.3 37.0
+MedAlpaca-7B Drugs 20.3 22.8 / 24.5 22.8 20.5 24.9 34.2 / 27.5 42.6 36.8
+Diseases 25.2 30.1 / 21.6 26.5 25.2 31.2 41.7 / 25.0 40.5 38.1
+Palmyra-Med-20B Drugs 30.1 31.4 / 31.3 29.1 29.6 28.5 37.3 / 31.9 35.4 35.0
+Diseases 33.2 35.1 / 30.2 32.3 34.5 41.1 48.4 / 39.6 46.3 50.1
+Table 2: Multi-query (here budget = 8) attack success rates (ASR %) on various LLMs by perturbing the drug and disease entity mentions in MedQA-USMLE and MedMCQA datasets.
+The PDWS attack uses the embedding from CODER and DiscreteZOO (DZOO) attacks use either CODER or GTE-base to calculate the cosine distance.
+For each model, the highest ASR for a type of entity-mention question in each dataset is bolded.
+5.2 Scaling characteristics from adversarial entities
+Size is not all for scaling model robustness.
+We compare model robustness using single-query ASR as the metric in Table 5 and Fig.
 3.
-$\frac{\Delta\text{Acc}}{\text{Acc}}=\frac{\Delta(n_{\text{TP}}+n_{\text{TN}})}{n_{\text{TP}}+n_{\text{TN}}}$ (6)
-R-CAP is equivalently the relative change in the number of correct model predictions (numbers of true positives $n_{\text{TP}}$ and true negatives $n_{\text{TN}}$) (Wang et al., 2023b).
-Tables 1-2 show that zero-shot adversarial sampling already incurs about 10-20% of relative change in accuracy.
-For the Flan-T5 series (Chung et al., 2022), R-CAP is more pronounced for smaller and less performant models, indicating the improvement of robustness as models scale up.
-However, the BLMs are generally more robust to entity substitution than GLMs for models of a similar size.
-The trend is consistent in both of the two regimes found through hyperparameter tuning of the powerscaled DWS.
-It is worth noting that although Palmyra-Med-20B (Kamble and Alshikh, 2023) has about three times the number of parameters than MedAlpaca-7B (Han et al., 2023), it is much more sensitive to entity perturbation.
-Therefore, the smaller BLM is more robust against perturbation.
-Another observation is that the R-CAP for models upon adversarial perturbations is asymmetrical.
-For most of the models, sampling adversarial entities closer to the key induces more performance drop than sampling further away.
+For the Flan-T5 series of GLMs (Chung et al., 2024) and Flan-UL2 (Tay et al., 2022), the single-query ASR is more pronounced in smaller and less performant models, indicating the improvement of robustness as the model’s size grows.
+However, an opposite trend of robustness is seem in BLMs.
+It is worth noting that although Palmyra-Med-20B (Kamble & Alshikh, 2023) has about three times the number of parameters than MedAlpaca-7B (Han et al., 2023), the larger model is much more sensitive to entity perturbation.
+In terms of ASR, the BLMs are comparable to the GLMs in each type of questions evaluated on.
+Another observation is that the ASR for models is asymmetrical under PDWS attacks.
+For most of the models, sampling adversarial entities with inverse-distance-proportional weights ($n<0$ in Eq.
+3) induces more performance drop than sampling with distance-proportional weights ($n>0$ in Eq.
+3), yet prominent counterexamples exist.
+The attack performance is further investigated in the sequential, multi-query setting (budget = 8) in Table 5.1, where we also compared sampling- against search-based DiscreteZOO attacks.
+Here, PDWS attacks in the $n<0$ regime are still largely the most effective.
+Figure 4: Example scaling curves of the query budget against ASR for disease-mention questions with (a) Flan-T5-xl on MedMCQA dataset and (b) MedAlpaca model on MedQA-USMLE dataset.
+The curves are generated using TCES of disease names with the methods described in the legends.
+The DiscreteZOO attacks were run in the low-query-budget setting.
+Executing the random sampling (RS) attack doesn’t require a text embedding, while the other attacks were evaluated with the CODER or GTE-base as text embedding.
+Sampling in specialized text embedding has favorable query scaling over blackbox gradient attacks.
+A query scaling study of adversarial attacks investigates attack success with varying query budgets per input instance (Shukla et al., 2021).
+It is a comprehensive evaluation of attack methods, but not yet frequently adopted in most publications, partly due to the computational cost.
+Fig.
+4 compares the query scaling of both sampling- and search-based methods for TCES.
+For low-query-budget attacks (Fig.
+1a), each scaling curve exhibits a rapid rise phase and a plateau phase, separated by an inflection point.
+The key message here is that the advantage of PDWS attacks over DiscreteZOO exists with a choice of text embedding.
+It is greater in specialized text embeddings but is reduced or diminished in general ones thanks to the difference in neighborhood structures.
+We expand the analysis in the following in three directions.
+  • Query scaling depends on text embedding: Changing the text embedding affects the query scaling behavior of both PDWS and DiscreteZOO attacks.
+Fig.
+4 shows that the influence of the embedding choice on PDWS is more pronounced in the rapid rise phase, while it influences the DiscreteZOO primarily in the plateau phase.
+Specifically, changing the text embedding from domain-specialized CODER (Yuan et al., 2022) to the general GTE-base (Li et al., 2023b), the performance advantage of PDWS over DiscreteZOO attacks shrunk in Flan-T5-xl (Fig.
+4a) and inverted in MedAlpaca-7B (Fig.
+4b).
+  • Sampling hyperparameter controls scaling curve shape in PDWS attacks: The hyperparameter $n$ changes the inflection point of the scaling curve.
+When the query budget is past the inflection point of the curve, random sampling starts to surpass PDWS ($n>0$) in ASR.
+This is a consequence of the distribution of adversarial entities in the embedding space and the regional bias derived from the sampling weights.
+The examples in Fig.
+4 show that the largest performance gap between the PDWS and DiscreteZOO attacks appears at around the inflection point of the respective query scaling curve.
+  • Gradient quality affects query scaling: The query complexity in blackbox gradient estimation is a long-standing theoretical topic in derivative-free optimization (L’Ecuyer & Yin, 1998) and online learning (Duchi et al., 2015).
+Nevertheless, in application settings such as adversarial machine learning, how gradient quality affects query scaling remains less studied for text data.
+Gradient estimation consumes a part of the total query budget, leading to a tradeoff between the gradient quality and terminal ASR attainable in the query scaling curve.
+We empirically explored the query scaling behavior of the DiscreteZOO attack in Appendix E with different candidate points in the multi-point gradient estimation, which in effect alters the gradient quality.
+The results show that having too few candidate points (i.e.
+small $M$ in Eq.
+6) leads to an early onset of the plateau phase, thereby limiting the highest attainable ASR within the total query budget, whereas having many candidate points can lead to a slower growth of ASR in the rapid rise phase of the query scaling curve.
 A 73-year-old man presents to the outpatient clinic complaining of chest pain with exertion.
@@
 (c) Adversarially perturbed
-Figure 4: (Top) An entity subsitution attack using n-acetylglucosamine as the replacement entity for metoprolol creates an adversarial distractor.
+Figure 5: (Top) An entity substitution attack using n-acetylglucosamine as the replacement entity for metoprolol creates an adversarial distractor.
 (Bottom) Heatmaps of token-wise Shapley values for a question before and after the adversarial attack on option C.
 The model prediction changes from the correct choice of B (unperturbed) to the incorrect choice of D (adversarially perturbed).
-Entity substitution attacks manipulate explainability.
+5.3 Adversarial entities manipulate explainability
 Post-hoc analysis is a common way to understand the rationale behind model decisions (Murdoch et al., 2019).
-Score-based feature importance such as that based on the Shapley values are routinely used to construct post-hoc explanation of model prediction through feature attribution (Chen et al., 2023).
-The stability of model explanation is a growing concern for their proper use (Alvarez-Melis and Jaakkola, 2018; Hancox-Li, 2020; Lin et al., 2023).
+Score-based feature importance such as that based on the Shapley values is routinely used to construct post-hoc explanation of model prediction through feature attribution (Chen et al., 2023).
+The stability of model explanation is a growing concern for their proper use (Alvarez-Melis & Jaakkola, 2018; Hancox-Li, 2020; Lin et al., 2023).
 Investigations on the susceptibility of LLM explanations to adversarial perturbation in QA are still lacking.
@@
 5 produced using Flan-T5-large.
-This behavior is an indicator of the symbol-binding characteristic of LLMs in multiple-choice QA (Robinson and Wingate, 2023).
+This behavior is an indicator of the symbol-binding characteristic of LLMs in multiple-choice QA (Robinson & Wingate, 2023).
 Moreover, the Shapley profile typically exhibits pronounced changes before and after entity perturbation, especially in the text locations close to the perturbed entities (Fig.
 5 and Appendix E).
-Together, these observations indicate that the substitution attacks effectively manipulate model explanations, because the explanation after attack aligns with the wrong prediction, as if they were correct.
-The behavior occurs despite unchanged context, stem, and key ($\mathcal{C}$, $\mathcal{S}$, and $k$), which should contain the most important information for understanding and answering the question (Chai and Jin, 2004).
-6 Related works
-Text substitution attack
-Adversarial attacks in neural language models have been demonstrated from the character to the sentence level (Zhang et al., 2020).
-The majority of text substitution attacks employ typos (Gao et al., 2018) and synonyms (Mozes et al., 2021), which effectively impose lexical constraints.
-Word frequency has been used as a metric to select adversarial examples in (Mozes et al., 2021).
-Synonym substitution attacks have been demonstrated in sentiment analysis (Liu et al., 2022).
-Alzantot et al.
-(2018) used a genetic algorithm to search for adversarial word substitution.
-Model sensitivity to entity-level perturbations such as substitution and swapping create problematic situations for knowledge-intensive tasks, such as knowledge conflicts (Longpre et al., 2021), compromised performance in machine reading comprehension (Yan et al., 2022) and table interpretation tasks (Koleva et al., 2023).
-Robustness in QA
-The robustness of QA systems has been studied using perturbative methods (Jia and Liang, 2017) and meaning-preserving transformations (Gan and Ng, 2019; Elazar et al., 2021).
-Sen and Saffari (2020) introduced a set of perturbations to assess language model generalization between QA datasets.
-Richardson and Sabharwal (2020) proposed a distractor design method using proximity in semantic space to probe model understanding of word senses.
-Awadalla et al.
-(2022) examined the connection between in- and out-of-distribution robustness for different LLMs.
-The performance of LLMs is sensitive to the ordering of options (Pezeshkpour and Hruschka, 2023; Zheng et al., 2023).
-Overall, the brittleness of knowledge in LLMs remains a primary limitation for their applications (Razniewski et al., 2021; Elazar et al., 2021; Ji et al., 2023).
-Adversarial attack characteristics
-The defining characteristic of adversarial attacks is to induce false predictions in trained models.
-Besides, attacks may possess transferability, universality (Zou et al., 2023), and they may obstruct model explanation (Noppel and Wressnegger, 2024).
-Transferability refers to the effectiveness of the same example being adversarial to a number of distinct models (Papernot et al., 2016).
-Universal adversarial perturbations are reusable across many examples tested on the same model (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a).
-Model explanations may be manipulated by specifically designed perturbations to the inputs (Dombrowski et al., 2019) or biased data subsampling (Laberge et al., 2023).
+Together, these observations indicate that the substitution attacks effectively manipulate model explanations, because the explanation after the attack aligns with the wrong prediction, as if they were correct.
+The behavior occurs despite unchanged context, stem, and key ($\mathcal{C}$, $\mathcal{S}$, and $k$), which should contain the most important information for understanding and answering the question (Chai & Jin, 2004).
 6 Discussion and future work
-The present study illustrates that side information introduced through entity substitution leads to noticeable degradation of model performance, by modifying only a single sampling step and without gradient computation.
-Moreover, type-constrained entity substitution makes sure that the replacement entity appears natural to human and is less likely to be detected using grammar check or topical filtering (Morris et al., 2020).
+The present study illustrates that distracting information introduced through QA distractors can lead to noticeable performance degradation in billion-parameter LLMs.
+TCES with only a single entity modification step combined with adversarial sampling methods and appropriate text embeddings is a feasible way for collecting adversarial entity data.
+Moreover, TCES ensures that the replacement entity appears natural to humans and is less likely to be detected using grammar check or topical filtering (Morris et al., 2020a).
 The setting simulates realistic scenarios where the “attacks" may be initiated by unsuspecting users, such as biomedical professionals using an LLM-based clinical decision support system (Liu et al., 2023).
 In the examples illustrated in Fig.
-4 and Appendix E, adversarial distractors can lead to disease misdiagnosis or prescription of the wrong medication, which are detrimental scenarios that may occur in algorithmic decision-making in biomedicine.
-Further improvement on the attack performance may be achieved by sequential sampling, an intrinsic approach to boost the sampling performance, or by changing the entity vocabulary of the perturbation set, an extrinsic approach that depends on the user and application setting.
+5 and Appendix E, adversarial distractors can lead to disease misdiagnosis or misprescription of medication, which are detrimental scenarios that may occur in algorithmic decision-making in biomedicine using LLMs.
+Further improvement on the attack performance may be achieved using learned samplers through rejection mechanisms (Narasimhan et al., 2024) or by finetuning the text embedding to improve the query efficiency of sampling.
 Our approach may be integrated into interactive platforms for adversarial testing in open-ended human-machine conversation (Chao et al., 2023) for online model auditing.
-Future research may focus on how to make LLMs more reliable through prompt engineering (Si et al., 2023; Nori et al., 2023) or leveraging retrieval from nonparametric knowledge bases (Weikum et al., 2021; Soman et al., 2023) or text sources (Wang et al., 2023c) to improve model generalization to long-tail entities.
-The DWS methodology may be extended to kernel methods (Hofmann et al., 2008) or metric learning (Bellet et al., 2015), which can capture more general distance dependence.
+Future research may focus on how to make LLMs more reliable through prompt engineering (Si et al., 2023; Nori et al., 2023) or leveraging retrieval from nonparametric knowledge bases (Weikum et al., 2021; Soman et al., 2023) or text sources (Wang et al., 2023b) to improve model generalization to long-tail entities.
 Besides vulnerability identification, the current work motivates adversarial defense strategies based on generalized distances to establish a tiered system for validating adversarial robustness in user-facing and domain-oriented applications.
-Limitations
-Although NEs convey parts of the key information in the question, they fall short of capturing the entirety of the textual meaning.
-To explore other attack surfaces within the adversarial landscape of QA, one would need to evaluate the robustness in multidimensional perturbation space (Wang et al., 2022).
-Besides, the meaning of Shapley values calculated here requires additional interpretation due to feature correlation.
-Although we use them in part to reveal the importance of specific tokens, their robustness is also worth exploring using sentence-level features (Chen et al., 2023).
-Ethics statement
-The scenarios considered in this work using public datasets relate to potential real-world deployment of LLMs in biomedical education and clinical decision support.
-Our results show that the lack of adversarial testing methods in high-stakes domains may incur serious consequences because of naturally looking adversarial examples.
+Broader Impact Statement
+The scenarios considered in this work using public datasets relate to the potential real-world deployment of LLMs in biomedical education and clinical decision support.
+Our results show that the lack of robustness assessment or adversarial testing methods in high-stakes domains may incur serious consequences because of naturally looking adversarial examples.
 The design and deployment of adversarial testing should be specific to and carried out at every stage of the model development cycle, paired with tools such as adversarially robust model explainability methods for understanding model behavior.
-Appendix A Adversarial distractor generation
+Appendix A Adversarial distractor generation (General case)
 We formulate here the general case of adversarial distractor generation when the choices contain text other than entities and one or more entities are present in a single choice.
@@
 If $d_{i,j}$ in the question is adversarially perturbed, the model elicits a false response, while it answers correctly to the unperturbed question $\mathcal{Q}$,
-$\mathrm{argmax}$ $\Pr(Y=\mathcal{K}|\mathcal{Q}(d_{i,j}\rightarrow d_{i,j}^{\prime}))$
-$\neq\mathrm{argmax}\Pr(Y=\mathcal{K}|\mathcal{Q}).$ (8)
+$\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=\mathcal{K}|\mathcal{Q}(d_{i,j}\rightarrow d_{i,j}^{\prime}))\neq\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=\mathcal{K}|\mathcal{Q}).$ (9)
 The general case is reduced to the special case described in Section 3 when $l_{i}=1$, $\mathcal{D}_{i}=d_{i,1}\coloneqq d_{i}$, and $\mathcal{K}=k$, which recover $\mathcal{D}=\{d_{i}\}_{i=1}^{l}$, and $\mathcal{Q}=(\mathcal{C},\mathcal{S},\mathcal{D},k)$.
@@
 Here we present a simple proof of the deterministic limit of the powerscaled distance-weighted sampling described in Eq.
-(3) of the main text.
+3 of the main text.
 Although the distance function $h$ is not given a specific form, we assume it to satisfy the basic properties of a metric.
@@
 B.3 The case of $n=0$
-When $n=0$, it is straightforward to show that $p_{j}|_{n=0}=1/m$, regardless of $j$ using the definition in Eq. (3). This scenario therefore corresponds to uniform random sampling.
-Answer the following question without explanation.
+When $n=0$, it is straightforward to show that $p_{j}|_{n=0}=1/m$, regardless of $j$ using the definition in Eq. 3. This scenario therefore corresponds to uniform random sampling.
+Prompt format Answer the following question without explanation.
 [[[[Content]]]]: A 23-year-old pregnant woman at 22 weeks gestation presents with burning upon urination.
@@
 Appendix D Prompt formatting
-We structured the prompt with bracketed tags to divide different parts (Context, Question, Answer).
+We structured the prompt with bracketed tags to divide different parts ([Context], [Question], [Answer]).
 An example taken from the MedQA-USMLE dataset (Jin et al., 2021) is shown below.
-For the MedMCQA dataset (Pal et al., 2022), only the Context and Answer [^1] Using the Context tag improves model performance than the Question tag here.
+For the MedMCQA dataset (Pal et al., 2022), only the [Context] and [Answer] [^1] Using the [Context] tag improves model performance than the [Question] tag here.
 tags were used because the questions are short and generally contain no context.
@@
 Appendix E Model evaluation and attack examples
+Model inference runtime
+The evaluations were carried out on a variety of NVIDIA GPUs depending on the model size.
+For smaller models such as the Flan-T5 series (Chung et al., 2024), the runtime on each type of question (involving drug or disease mentions) in a dataset was generally between 10 and 30 minutes, while it was from one to more than four hours for larger model such as Flan-UL2 (Tay et al., 2022) and Palmyra-Med-20B (Kamble & Alshikh, 2023).
+The runtime for each type of questions is between two and four hours.
+Figure 6: Quality of blackbox gradient estimation affects the scaling behavior of DiscreteZOO (DZOO) attacks using the CODER and GTE-base text embeddings.
+Multi-point gradient estimation
+We demonstrate the effects of the quality of gradient estimation on the performance of the blackbox gradient attack (DiscreteZOO (Berger et al., 2021)) using a representative dataset in Fig.
+6 using MedAlpaca-7B (Han et al., 2023) and the MedMCQA dataset (Pal et al., 2022).
+The evaluation used TCES introduced in Algorithm 1.
+According to Eq.
+6, for each entity to attack, we randomly sampled two, five, and ten candidate points in its neighborhood in the embedding space for gradient estimation.
+The attack on each input instance proceeds with a fixed total of LLM queries (total query budget) as indicated on the horizontal axis.
+The results show that the larger the number of candidate points used for gradient estimation, the more effective the attack becomes before it consumes the total query budget, as measured by the attack success rate (ASR).
+Having too few candidate points can result in an early plateau phase in the ASR, which stunts its growth.
+Yet having more candidate points can slightly reduce the ASR growth in the rapid rise phase of the scaling curve.
+The general trends are consistent in both the CODER (Yuan et al., 2022) and GTE-base (Li et al., 2023b) embeddings used for the illustration.
+Perturbed explaination in adversarial examples
 We list some examples of successful entity substitution attacks using adversarial distractors following Fig.
@@
 Beneath the text box are the corresponding heatmaps of token-wise Shapley values (or Shapley profile) before and after the entity substitution attack.
-The Shapley profiles for the MedQA-USMLE dataset (Jin et al., 2021) were generated using the model Flan-T5-large and those for the MedMCQA (Pal et al., 2022) using the model Flan-T5-xxl (Chung et al., 2022).
+The Shapley profiles for the MedQA-USMLE dataset (Jin et al., 2021) were generated using the model Flan-T5-large and those for the MedMCQA (Pal et al., 2022) using the model Flan-T5-xxl (Chung et al., 2024).
 The Shapley values were calculated using the partition explainer in shap, which uses the Owen values as an approximation (Chen et al., 2023).
-Of the LLMs evaluated in this work, the proprietary models (GPT-3.5, PaLM2) were accessed through third-party APIs, while the open-source models were accessed on local cache to improve speed.
-The evaluations were carried out on a variety of NVIDIA GPUs depending on the model size.
-For smaller model such as the Flan-T5 series (Chung et al., 2022), the runtime on each type of questions (involving drug or disease mentions) in a dataset was generally between 10 and 30 minutes, while it was from one to more than four hours for larger model such Flan-UL2 (Tay et al., 2022) and Palmyra-Med-20B (Kamble and Alshikh, 2023).
-For proprietary models accessed through their APIs, the runtime for each type of questions is between two and four hours.
-An 18-year-old man presents with a sudden loss of consciousness while playing college football.
+MedQA-USMLE An 18-year-old man presents with a sudden loss of consciousness while playing college football.
 There was no history of a concussion.
@@
 (c) Adversarially perturbed
-A 41-year-old woman comes to the physician because of a 3-month history of anxiety, difficulty falling asleep, heat intolerance, and a 6-kg (13.2-lb) weight loss.
+MedQA-USMLE A 41-year-old woman comes to the physician because of a 3-month history of anxiety, difficulty falling asleep, heat intolerance, and a 6-kg (13.2-lb) weight loss.
 The patient’s nephew, who is studying medicine, mentioned that her symptoms might be caused by a condition that is due to somatic activating mutations of the genes for the TSH receptor.
@@
 (g) Adversarially perturbed
-A 25-year-old man is brought to the emergency department after his girlfriend discovered him at home in a minimally responsive state.
+MedQA-USMLE A 25-year-old man is brought to the emergency department after his girlfriend discovered him at home in a minimally responsive state.
 He has a history of drinking alcohol excessively and using illicit drugs.
@@
 (k) Adversarially perturbed
-A 70-year-old woman presents to the office for a yearly physical.
+MedQA-USMLE A 70-year-old woman presents to the office for a yearly physical.
 She states she has recently started experiencing pain in her legs and her back.
@@
 (o) Adversarially perturbed
-Which of the following is granted orphan drug status for the treatment of Dravet’s syndrome?
+MedMCQA Which of the following is granted orphan drug status for the treatment of Dravet’s syndrome?
 A: Stiripentol A: Stiripentol B: Icatibant B: Tocilizumab   ✓ C: Pre-vitamin D3 C: Pitolisant D: Tafamidis   ✓ D: Tafamidis
@@
 (s) Adversarially perturbed
-Which among the following may be used as a sedatives-hypnotic? A: Zolmitriptan A: Zolmitriptan   ✓ B: Zileuton B: Zileuton C: Zolpidem   ✓ C: N-acetylglucosamine D: Zalcitabine D: Zalcitabine
+MedMCQA Which among the following may be used as a sedatives-hypnotic? A: Zolmitriptan A: Zolmitriptan   ✓ B: Zileuton B: Zileuton C: Zolpidem   ✓ C: N-acetylglucosamine D: Zalcitabine D: Zalcitabine
 (v) Unperturbed

### v3

@@
 adversarial entities) in natural language processing tasks raises questions about their potential impact on the knowledge robustness of pre-trained and finetuned LLMs in high-stakes and specialized domains.
-We examined the use of type-consistent entity substitution as a template for collecting adversarial entities for billion-parameter LLMs with biomedical knowledge.
-To this end, we developed an embedding-space attack based on powerscaled distance-weighted sampling to assess the robustness of their biomedical knowledge with a low query budget and controllable coverage.
-Our method has favorable query efficiency and scaling over alternative approaches based on random sampling and blackbox gradient-guided search, which we demonstrated for adversarial distractor generation in biomedical question answering.
-Subsequent failure mode analysis uncovered two regimes of adversarial entities on the attack surface with distinct characteristics and we showed that entity substitution attacks can manipulate token-wise Shapley value explanations, which become deceptive in this setting.
-Our approach complements standard evaluations for high-capacity models and the results highlight the brittleness of domain knowledge in LLMs [^1] The code and datasets developed for the work is available at https://github.com/RealPolitiX/qstab/..
+We examined the use of type-consistent entity substitution as a template for collecting adversarial entities for medium-sized billion-parameter LLMs with biomedical knowledge.
+To this end, we developed an embedding-space, gradient-free attack based on powerscaled distance-weighted sampling to assess the robustness of their biomedical knowledge with a low query budget and controllable coverage.
+Our method has favorable query efficiency and scaling over alternative approaches based on blackbox gradient-guided search, which we demonstrated for adversarial distractor generation in biomedical question answering.
+Subsequent failure mode analysis uncovered two regimes of adversarial entities on the attack surface with distinct characteristics.
+We showed that entity substitution attacks can manipulate token-wise Shapley value explanations, which become deceptive in this setting.
+Our approach complements standard evaluations for high-capacity models and the results highlight the brittleness of domain knowledge in LLMs [^1] The code and datasets developed for the work are available at https://github.com/RealPolitiX/qstab..
 1 Introduction
 LLMs such as pre-trained and finetuned generalist language models (GLMs) and their domain-adapted versions (Zhao et al., 2023; Wang et al., 2023a) incorporating transformer architectures are emerging as the technological backbone of next-generation search engines and knowledge bases (Petroni et al., 2019; Sung et al., 2021).
-They are increasingly used in knowledge-intensive tasks by the general public without sufficient assessment of inherent issues in trustworthiness and safety (Anonymous, 2023; Spitale et al., 2023; Si et al., 2023).
+They are increasingly used in knowledge-intensive tasks by the general public without sufficient assessment of inherent issues in trustworthiness and safety (Clusmann et al., 2023; Spitale et al., 2023; Si et al., 2023).
 Adversarial examples are valuable probes for model vulnerability and robustness, and are essential for performance improvements and regulatory audits (Shayegani et al., 2023).
@@
 For language models, adversarial examples are generated through guided search (Alzantot et al., 2018), sampling and probabilistic methods (Ren et al., 2019; Yang et al., 2020; Yan et al., 2022), or collected dynamically by human operators (Jia & Liang, 2017; Wallace et al., 2019b).
-Most of these methods require a large amount of model evaluations (in the hundreds to thousands or beyond per instance) or whitebox access to model internals, and their evaluations tend to focus on small models (with < 1B parameters) (Maheshwary et al., 2021), which are insufficient in the era of LLMs.
+Most of these methods require a large amount of model evaluations (in the hundreds to thousands or beyond per instance) or whitebox access to model internals, and their evaluations tend to focus on small models (with < 1B parameters) (Maheshwary et al., 2021; Yu et al., 2024), which are insufficient in the era of LLMs.
 Sampling-based attacks obviate the need for predetermined search heuristics and can operate with only blackbox model access and a low query budget.
 Demonstrations of query-efficient attacks on LLMs meet an emerging demand for their continuous monitoring (Metaxa et al., 2021), which is particularly important for deploying LLMs in specialized domains.
-Question answering (QA) has been the proving ground in recent achievements of biomedical language models (BLMs), here referring to GLMs adapted to the biomedical domain (Singhal et al., 2023a; Kung et al., 2023; Singhal et al., 2023b).
+Question answering (QA) has been the proving ground in recent achievements of biomedical language models (BLMs), here referring to GLMs adapted to the biomedical domain (Singhal et al., 2023; Kung et al., 2023; Liévin et al., 2024; Saab et al., 2024).
 Knowledge and reasoning components in biomedical QA assess the understanding of diverse concepts, which are represented by domain-specific named entities (NEs).
-The present work focuses on the identification of adversarial entities in biomedical QA for both GLMs and BLMs.
-While GLMs have general biomedical knowledge during model development, BLMs gain further domain knowledge primarily through finetuning on a variety of text data from clinical notes (Lehman et al., 2023), electronic health records (Wornow et al., 2023), to abstracts or excerpts from scholarly publications (Tinn et al., 2023).
-Although QA is frequently employed paradigm for model performance evaluation (Gardner et al., 2019; Robinson & Wingate, 2023), currently, the performance evaluation in biomedical QA lacks robustness assessment and their deployment in the real world, albeit promising, is still in its infancy (Kung et al., 2023).
-Analyzing biomedical texts using LLMs requires them to have robust contextual knowledge and language understanding of entities composed of single or multiple words.
-Evaluating model robustness in prediction for high-stakes, specialized domains requires customized methods apart from traditional benchmarks.
+The present work aims to identify adversarial entities in biomedical QA for both GLMs and BLMs.
+While GLMs may have acquired general biomedical knowledge during model development, BLMs gain further domain knowledge primarily through finetuning on a variety of text data from clinical notes (Lehman et al., 2023), electronic health records (Wornow et al., 2023), to abstracts or excerpts of scholarly publications (Tinn et al., 2023).
+Although QA is a frequently employed paradigm for model performance evaluation (Gardner et al., 2019; Robinson & Wingate, 2023), currently, biomedical QA still lacks domain-specific robustness assessment and its use in the real-world applications, albeit promising, is still in its infancy (Kung et al., 2023).
+Evaluating model robustness for high-stakes, specialized domains requires customized perturbations apart from traditional benchmarks (Cecchini et al., 2024).
+Deployable LLMs in the biomedical domain should have robust contextual knowledge and language understanding of domain-specific entities.
 (a) (a)
@@
 (c) (b)
-Figure 1: Text substitution attack on QA with adversarial distractors.
+Figure 1: Entity substitution attack on QA with adversarial distractors.
 (a) Typical query scaling curves in low- and high-query-budget attack settings.
@@
 $\mathcal{E}$ represents the vocabulary set.
-D is the key to the question, A-C the original distractors, C is an adversarial distractor at distance $h$ from the center at key D.
-The locations of A-C and C are relative to D in the embedding space.
-In this work, we developed an effective in-domain, query-efficient adversarial probe for the knowledge robustness of LLMs in entity-rich domains using type-consistent entity substitution (TCES).
-They operate when a large number of viable replacements ($N_{\mathcal{E}}$ > 1000) per attack exist, while only a small number of text locations are permissible.
-The setting is prevalent in specialized domains where the NEs have shared characteristics such as their type and (word or phrase) structure.
-It differs from traditional synonym substitution attacks (Liu et al., 2022), which use a much smaller number of viable replacements per attack with matching word senses but can operate on many potential text locations.
-Our focus is on generating adversarial examples of distractors (see Fig.
+D is the key to the question, A-C the original distractors, C is an adversarial distractor at distance $h$ from the key D.
+In this work, we proposed and compared sampling- and search-based query-efficient adversarial attacks for evaluating the knowledge robustness of LLMs in the entity-rich, biomedical domain using type-consistent entity substitution (TCES).
+We investigated the setting where a large number of viable replacements ($N_{\mathcal{E}}$ > 1000) per attack exist at only a small number of permissible text locations, which is prevalent in specialized domains where the NEs have shared characteristics such as their type and (word or phrase) structure.
+The setting differs from traditional synonym substitution (Liu et al., 2022; Yu et al., 2024), which have a significantly smaller number of viable replacements with matching word senses but can operate on many potential text locations.
+To relate to real-world scenarios (Apruzzese et al., 2023), we consider the attacks in the blackbox and hard-label setting with a limited query budget ($B\ll N_{\mathcal{E}}$).
+We used the attacks to generate adversarial examples of distractors (see Fig.
 1b), or adversarial distractors in biomedical QA tasks.
-For relevance to real-world scenarios (Apruzzese et al., 2023), we consider the attacks in the blackbox and hard-label setting with a limited query budget ($B\ll N_{\mathcal{E}}$).
 Our main contributions include: (i) Introduction of powerscaled distance-weighted sampling (PDWS) method in embedding space to generate adversarial entities.
 Our work reports the first use of DWS for adversarial attacks in text [^2] Throughout the text, we use DWS to refer to the class of sampling method that employs distance-dependent probability weights.
-The powerscaled DWS is a type of DWS, so is the other version introduced in (Wu et al., 2017)..
+The powerscaled DWS is a type of DWS, so is the other version introduced in Wu et al.
+(2017)..
 The attack scheme has a broad coverage of the embedding space centered around an anchor point (illustrated as the key D in Fig.
-1c) and maps the entity-level robustness landscapes of different LLMs in two regimes.
+1c) and maps the entity-level robustness landscapes of different LLMs in two regimes (near and far).
 (ii) Introduction of the FDA-drugs and CTD-diseases datasets as the source vocabularies of the perturbation sets, along with an entity type-annotated version of existing biomedical QA datasets MedQA-USMLE (Jin et al., 2021) and MedMCQA (Pal et al., 2022).
@@
 1b).
-(iii) We unified sampling- and gradient search-based methods in TCES as an attack template, which we used for quantitative comparison of the scaling behaviors of the attacks in the same footing.
-We found that in the blackbox setting, sampling attacks can have favorable query scaling over gradient-based attacks, when are limited by the quality of gradient estimation within a fixed query budget.
+(iii) We unified sampling- and gradient search-based methods in discrete text space and formulated TCES as an attack template, which we used for quantitative comparison of the scaling behaviors of the attacks in the same footing.
+We found that in the blackbox setting, sampling attacks can have favorable query scaling over gradient-based attacks, which are gradient quality-limited within a fixed query budget.
 (iv) By investigating the relationship between Shapley value-based model explanation (Chen et al., 2023) and adversarial attacks, we uncovered the defining signatures of successful attacks as a result of entity substitution.
 2 Related works
-Text substitution attack
+Text substitution attacks
 Adversarial attacks in neural language models have been demonstrated from the character to the sentence level (Zhang et al., 2020).
@@
 Word frequency has been used as a metric to select adversarial examples in (Mozes et al., 2021).
-Synonym substitution attacks have been demonstrated in sentiment analysis (Liu et al., 2022).
+Synonym substitution attacks have been demonstrated in sentiment analysis (Liu et al., 2022; Yu et al., 2024).
 Alzantot et al.
 (2018) used a genetic algorithm to search for adversarial word substitution.
-Model sensitivity to entity-level perturbations such as substitution and swapping create problematic situations for knowledge-intensive tasks, including knowledge conflicts (Longpre et al., 2021), compromised performance in machine reading comprehension (Yan et al., 2022) and table interpretation tasks (Koleva et al., 2023).
+Model sensitivity to entity-level perturbations including substitution and swapping create problematic situations such as knowledge conflicts (Longpre et al., 2021) for knowledge-intensive tasks, compromised performance in machine reading comprehension (Yan et al., 2022) and table interpretation tasks (Koleva et al., 2023).
 Robustness in QA
@@
 (2022) examined the connection between in- and out-of-distribution robustness for different LLMs.
-The performance of LLMs is sensitive to the ordering of options (Pezeshkpour & Hruschka, 2023; Zheng et al., 2024).
+The performance of LLMs is sensitive to the ordering of choices (Pezeshkpour & Hruschka, 2023; Zheng et al., 2024).
 Overall, the brittleness of knowledge in LLMs remains a primary limitation for their applications (Elazar et al., 2021; Augenstein et al., 2024).
-Adversarial attack characteristics
-The defining characteristic of adversarial attacks is to induce false predictions in trained models.
-Other characteristics include (i) their generality such as transferability, universality (Zou et al., 2023).
+Characteristics of adversarial examples
+The defining characteristic of adversarial examples is the induction of false predictions in trained models that would otherwise predict accurately.
+Other characteristics include (i) their generality such as transferability and universality (Zou et al., 2023).
 Transferability refers to the effectiveness of the same example being adversarial to distinct models (Papernot et al., 2016).
 Universal adversarial perturbations are reusable across many examples tested on the same model (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a).
-(ii) The quality of adversarial examples may be judged by their perceptibility to human (Dyrmishi et al., 2023).
-(iii) Adversarial attacks can operate in the low- and high-query-budget settings (Shukla et al., 2021).
-(iv) Attacked text using specifically designed perturbations to the inputs can obstruct model explanation (Noppel & Wressnegger, 2024) and render it unuseful.
-3 Entity-centric adversarial attack
+(ii) The quality of adversarial examples may be judged by their perceptibility to human (Dyrmishi et al., 2023) or simple defense mechanisms such as grammar check.
+(iii) Adversarial attacks can operate in the low- and high-query-budget settings (Shukla et al., 2021) with different query scaling behaviors.
+(iv) Adversarial examples by specifically designed perturbations can obstruct model explanation (Noppel & Wressnegger, 2024) and render it unuseful.
+3 Entity-centric adversarial attacks
 3.1 Adversarial distractor generation
-We focus on constructing adversarial distractors in multiple-choice QA.
-Let each question $\mathcal{Q}$ be represented by a tuple, $\mathcal{Q}=(\mathcal{C},\mathcal{S},\mathcal{D},k)$, with the context $\mathcal{C}$, question stem $\mathcal{S}$ [^3] The stem in the example of Fig.
-1b is the last sentence starting with “which” on the left side.
-The context is the text before the stem., a distractor set $\mathcal{D}=\{d_{i}\}_{i=1}^{l}$, and a key $k$.
+Definition 3.1.
+A multiple-choice question $\mathcal{Q}$ is represented by a tuple, $\mathcal{Q}=(\mathcal{C},\mathcal{S},\mathcal{D},k)$, with the context $\mathcal{C}$, question stem $\mathcal{S}$, a distractor set $\mathcal{D}=\{d_{i}\}_{i=1}^{l}$, and a key $k$.
 The distractors are the wrong choices given to confuse test-takers, while the key is the correct one (CH & Saha, 2020).
-QA is a natural language generation task.
-In an autoregressive LLM, the answer corresponds to the next tokens generated following $(\mathcal{C},\mathcal{S},\mathcal{D},k)$.
-The adversarial distractor is a threat model in which the perturbed distractor set $\mathcal{D}^{\prime}$ takes the form
-$\mathcal{D}^{\prime}=(\mathcal{D}\setminus\{d\})\cup\{d^{\prime}\},$ (1)
+Definition 3.2 (Adversarial distractor).
+Given the original (or unperturbed) and a perturbed multiple-choice question pair, $\mathcal{Q}=(\mathcal{C},\mathcal{S},\mathcal{D},k)$ and $\mathcal{Q}^{\prime}=(\mathcal{C},\mathcal{S},\mathcal{D}^{\prime},k)$, which differ by only a distractor, $\mathcal{D}^{\prime}\setminus(\mathcal{D}\cap\mathcal{D}^{\prime})=\{d^{\prime}\}$.
+The distractor $d^{\prime}$ is adversarial if the perturbed question elicits a false response from the model, while it answers correctly to the unperturbed question,
+$\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=k|\mathcal{C},\mathcal{S},\mathcal{D}^{\prime},k)\neq\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=k|\mathcal{C},\mathcal{S},\mathcal{D},k).$ (1)
 […]
-$\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=k|\mathcal{C},\mathcal{S},\mathcal{D}^{\prime},k)\neq\mathrm{argmax}\,{\Pr}_{\mathrm{LLM}}(Y=k|\mathcal{C},\mathcal{S},\mathcal{D},k).$ (2)
-The formulation in Eq.
-2 is an untargeted attack, while the targeted version only requires restricting the answer to the perturbed question to a specific distractor.
-We extend Eq.
-2 to a more general case in Appendix A with multiple entities and additional text in the distractors.
+Example 3.1.
+In the example of Fig.
+1b, the stem $\mathcal{S}$ is the last sentence on the left side, “Which of the following is most likely to be elevated in this patient?” The context $\mathcal{C}$ is all the text before the stem, “A 50-year-old woman … Her pupils are dilated.” The original distractor set $\mathcal{D}$ = {Serum creatinine, Temperature, Creatine phosphokinase, Blood pressure}, the perturbed distractor set $\mathcal{D}^{\prime}$ = {Serum creatinine, Temperature, Parathyroid hormone, Blood pressure}, the key $k$ is Blood pressure, or equivalently, D [^3] For simplicity, in this work, we don’t use different symbols to distinguish between the content of a distractor (or a key) from its index (e.g.
+A, B, C, D)..
 Unlike text addition attacks that introduce distracting information (Jia & Liang, 2017; Wallace et al., 2019a), the edit constraint on $\mathcal{Q}$ that is usually enforced in text adversarial attacks (Roth et al., 2024) doesn’t preclude large semantic perturbations at the entity level in the distractors.
@@
 In practice, $\mathcal{E}$ may be compiled from potential model use cases or curated by regulatory bodies in a model audit.
+Next, we discuss a unified view of discovering adversarial entities in the embedding space using blackbox (or gradient-free) attack methods based on sampling and search.
 3.2 Powerscaled distance-weighted sampling
+Definition 3.3 (Text span representations).
+Given the token sequence representation of a span $e$, we define its corresponding vector representation in a text embedding of dimension $l$ as e $\in\mathbb{R}^{l}$.
 We consider sampling with distance-dependent probability weights from the embedding space of NEs.
@@
 We propose a more flexible version independent of the embedding dimensionality using a powerscaled distance with a tunable hyperparameter.
-Specifically, for a key $k$, we assign a distance-dependent probability weight $p_{j}$ to entity $e_{j}$ such that the distractor $d$ is substituted by
+Specifically, for a key $k$, we assign a distance-dependent probability weight $p_{j}\in[0,1]$ to entity $e_{j}$ such that the distractor $d$ is substituted by
 $d^{\prime}=e_{j}\quad\mathrm{w.p.}\quad p_{j}=\frac{h^{n}(k,e_{j})}{\sum_{j}h^{n}(k,e_{j})}.$ (2)
@@
 When $n>0$, as $n$ increases, the entities further away from $k$ are more preferably weighted.
-When $n<0$, the sampling method employs inverse distance weights, and the entities closer to $k$ are more preferably weighted.
-In the subsequent evaluations, we use $h(k,e_{j})=\texttt{CosineDist}(\mathbf{k},\mathbf{e}_{j})=1-\mathbf{k}\cdot\mathbf{e}_{j}/(\|\mathbf{k}\|\cdot\|\mathbf{e}_{j}\|)$, calculated in a text embedding of choice.
-We use here the italic letters ($k,e$) to represent text while the bolded letters ($\mathbf{k},\mathbf{e}$) are their corresponding vector representations.
+When $n<0$, the sampling method employs inverse-distance-scaled weights, and the entities closer to $k$ are more preferably weighted.
+In the subsequent evaluations, we use $h(k,e_{j})=\texttt{CosineDist}(\mathbf{k},\mathbf{e}_{j})=1-\mathbf{k}\cdot\mathbf{e}_{j}/(\|\mathbf{k}\|\cdot\|\mathbf{e}_{j}\|)$, calculated in a text embedding of choice, where $\mathbf{k}$ and $\mathbf{e}$ are the respective vector representations of $k$ and $e$.
 The formalism accommodates a deterministic regime as a limiting case [^4] The limit $n\rightarrow-\infty$ indicates replacing the distractor $d$ with $k$ at all times.
-Since $k$ is not in the perturbation set -$\mathcal{E}_{\textrm{-}k}$, this limit is not attained..
-It corresponds to $n\rightarrow+\infty$, when $p_{j}$ is vanishingly small for all but the entity furthest from $k$,
+Since $k$ is not in the perturbation set -$\mathcal{E}_{\textrm{-}k}$, this limit is not attained., when $n\rightarrow+\infty$, and $p_{j}$ is vanishingly small for all but the entity furthest from $k$,
 $d^{\prime}=e_{j}\quad\mathrm{s.t.}\quad e_{j}=\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmax}}\,h(k,e_{j}).$ (3)
@@
 Therefore, our approach bridges the probabilistic and deterministic approaches to sampling adversarial examples.
-3.3 Sampling in zeroth-order gradient attack
-Adversarial attacks based on zeroth-order optimization (ZOO) use estimated gradients from finite difference to guide the search (Ilyas et al., 2018; Cheng et al., 2019).
+3.3 Sampling view of zeroth-order adversarial attack
+Blackbox attacks using approximate gradients may proceed with zeroth-order optimization (ZOO), where the gradients are estimated by finite difference to guide the search (Ilyas et al., 2018; Cheng et al., 2019).
 In the discrete text space, Berger et al.
 (2021) developed the DiscreteZOO attack, which contains three components: the word importance ranker, the candidate sampler, and the gradient-based optimizer.
-It operates on the word level and is validated on BERT-sized models for synonym substitution attacks.
-We discuss here the last two components, which, in each iteration, aim to find a replacement text span $e^{\prime}$ through the gradient update rule
-$\mathbf{e}^{\prime}=\mathbf{e}_{0}-\lambda\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0}),$ (5)
-where $e_{0}$ is the original text span, $\lambda$ the learning rate, and ^$\widehat{\nabla}_{\mathrm{ZO}}$ indicates the zeroth-order gradient estimator.
-We use $\mathrm{LLM}(\mathcal{Q};e_{0})$ to denote the LLM output with the unperturbed question $\mathcal{Q}$ (containing $e_{0}$) as the input, then the multi-point ($M\geq 2$) gradient estimate is constructed by querying the LLM $M$ times and computing
+It operates at the word level and has been validated on BERT-sized models for synonym substitution attacks.
+The last two components aim to find a replacement text span $e^{\prime}$ in each iteration through the gradient update rule
+$\mathbf{e}^{\prime}=\mathbf{e}_{0}-\lambda\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0}).$ (4)
+Here, $e_{0}$ refers to the original text span, $\mathbf{e}^{\prime}$ and $\mathbf{e}_{0}$ are the vector representations of $e^{\prime}$ and $e_{0}$, respectively.
+$\lambda$ the learning rate, ^$\widehat{\nabla}_{\mathrm{ZO}}$ indicates the zeroth-order gradient estimator, and ^$\widehat{\nabla}_{\mathrm{ZO}\mathrm{LLM}(\mathcal{Q};e_{0})}$ has the same dimensionality as $\mathbf{e}^{\prime}$ or $\mathbf{e}_{0}$.
+We use $\mathrm{LLM}(\mathcal{Q};e_{0})$ to denote the LLM text output with the unperturbed question $\mathcal{Q}$ (containing $e_{0}$) as the input, then the multi-point ($M\geq 2$) gradient estimate is constructed by querying the LLM $M$ times and computing
 ^$\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0})=\mathbb{E}_{\textbf{u}}\left[\frac{\Delta\mathrm{LLM}(\mathcal{Q};e_{0})}{\Delta\mathbf{e}_{0}}\textbf{u}\right]=\frac{1}{M}\sum_{i=1,e_{i}\in\mathcal{E}_{\textrm{-}k}}^{M}\frac{\left(\mathrm{LLM}(\mathcal{Q^{\prime}};e_{i})-\mathrm{LLM}(\mathcal{Q};e_{0})\right)\cdot(\mathbf{e}_{i}-\mathbf{e}_{0})}{\|\mathbf{e}_{i}-\mathbf{e}_{0}\|^{2}}.$ (5)
@@
 5 by random sampling of neighbors (in candidate sampler) within a similarity threshold for synonyms.
-Given a noisy gradient estimate ^$\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0})$ and the discrete nature of the search space, we can rewrite Eq.
+The procedure was iterated once for all ranked attack locations.
+Given a noisy gradient estimate ^$\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0})$ and a discrete search space, we can rewrite Eq.
 4 effectively as deterministic DWS in the embedding space by
@@
 Therefore, in the budgeted setting, a tradeoff should exist between the efficacy of the attack and the number of queries consumed for each gradient estimation.
-4 Implementing attacks on biomedical question answering
+4 Implementing attacks on biomedical QA
 4.1 Dataset selection
@@
 The MedMCQA dataset contains short questions used in medical exams in India.
-They have been used in recent works for assessing the biomedical knowledge in LLMs (Singhal et al., 2023a; Han et al., 2023; Singhal et al., 2023b).
+They have been used in recent works for assessing the biomedical knowledge in LLMs (Singhal et al., 2023; Han et al., 2023; Liévin et al., 2024; Saab et al., 2024).
 Both datasets are publicly available and don’t contain personal information.
 We annotated the NEs in the two datasets according to the entity types of the Unified Medical Language System (UMLS) (Bodenreider, 2004) using scispaCy (Neumann et al., 2019).
-We then divided the QA datasets into drug- or disease-mention questions by the binary mention labels constructed according to the UMLS entity types (see Appendix C).
-4.2 Type-consistent entity substitution
-Implementing attacks that target entities requires accounting for type consistency and compatibility with simultaneous multi-word (or span-level) operations, because of the inseparable nature of the entity components (e.g.
-growth hormone, acquired immunodeficiency syndrome).
-Algorithm 1 Type-consistent entity substitution (TCES) template for collecting adversarial entities in distractors.
-Inputs: Question Q, LLM, entity type $\tau$, query budget $B$.
-Internals: Number of choices $N_{\text{ch}}$, text t, entity Ent, embedding Emb, correct answer k, and evaluator function $\mu$.
-Outputs: Number of queries and the replacement entity, None if unsuccessful when the query budget is exhausted.
+We divided the QA datasets into drug- or disease-mention questions according to the UMLS entity types (see Appendix C) of NEs in distractors.
+Algorithm 1 TCES attack template for collecting adversarial entities in distractors.
+Inputs: Question Q, LLM, entity type $\tau$, budget $B$.
+Internals: Number of choices $N_{\text{ch}}$, text t, entity Ent, embedding Emb, key or correct answer k.
+Outputs: Number of queries and the replacement entity, None if unsuccessful after all attempts.
  TCESAttacker(Q, Model, $\tau$, $B$)
@@
      return
+4.2 Type-consistent entity substitution (TCES)
+Implementing attacks that target entities requires accounting for type consistency and compatibility with simultaneous multi-word (or span-level) operations, because of the inseparable nature of the entity components (e.g.
+growth hormone, acquired immunodeficiency syndrome).
 This effectively results in more substitution patterns (one to two words, two to one word, two to two words, two to three words, etc.) than the more common one-to-one synonym substitution (Mozes et al., 2021).
-We describe type-consistent entity substitution (TCES) in Algorithm 1 as the task-agnostic attack template.
-TCES is a general and contextualized attack for entities, with the goal of adversarially modifying model outputs with only one span-level substitution, implemented using token boundaries.
-The template can accommodate single or multi-query attacks and it is applicable to any task that may be formatted into QA (Gardner et al., 2019).
+We describe TCES in Algorithm 1 as a general and contextualized attack template for entities, with the goal of adversarially modifying model outputs with only one span-level substitution, specified by token boundaries.
+The template can accommodate single- or multi-query attacks and it is applicable to any task that may be formatted into QA (Gardner et al., 2019).
 […]
@@
 The criterion we adopted here for a model to have biomedical knowledge is that it should have a baseline performance better than random guessing (e.g.
-an accuracy of 0.25 for a specific four-choice QA task) in the relevant domain-specific task.
+an accuracy of 0.25 for a specific four-choice QA task) in the evaluated domain-specific task.
 For GLMs, we used instruction-finetuned T5 (Flan-T5) series of models (Chung et al., 2024) and UL2 model (Tay et al., 2022) (Flan-UL2).
-For BLMs, we used MedAlpaca-7B (Han et al., 2023), Llama2-Medtuned-13B (LM2-Medtuned-13B) (Rohanian et al., 2024) and Writer’s Palmyra-Med-20B (Kamble & Alshikh, 2023).
-The selected models contain diverse encoding schemes and architectures and their sizes are $\sim$ 1-20B.
+For BLMs, we used MedAlpaca-7B (Han et al., 2023), MedLlama-3-8B-v2.0 from John Snow Labs (JSL-MedLM3-8Bv2) [^7] https://huggingface.co/johnsnowlabs/JSL-MedLlama-3-8B-v2.0, Llama2-Medtuned-13B (LM2-Medtuned-13B) (Rohanian et al., 2024), and Palmyra-Med-20B (Kamble & Alshikh, 2023) from Writer.
+The selected open-source models have sizes in $\sim$ 1-20B range.
+Besides, we also evaluated GPT-3.5 from OpenAI.
 Attack settings
-To single out the effect of the attacks, all models were evaluated at zero temperature or in the non-sampling setting and model inference was conducted in the zero-shot setting with only basic prompt instructions (see the prompt structure in Appendix D).
+All models were evaluated at zero temperature or in the non-sampling setting and model inference was conducted in the zero-shot setting with only basic prompt instructions (see the prompt structure in Appendix D).
 Model inference of Palmyra-Med-20B (Kamble & Alshikh, 2023) used 4-bit quantization to improve speed (Dettmers & Zettlemoyer, 2023).
@@
 The DiscreteZOO attack (Berger et al., 2021) was originally implemented in the textattack framework (Morris et al., 2020b), which we updated to be compatible with span-level perturbation using its boundaries, multi-GPU split-model inference, and decoder-based billion-parameter LLMs by modifying the attack recipe.
-To demonstrate the generality of the attack settings, we focus on the following three aspects in the benchmarks.
-  • Query budget: We used a fixed query budget for three types of model evaluations: (i) Single-query (budget = 1) sampling-based attacks were used as the reference because the DiscreteZOO attack requires a minimum of 3 model queries; (ii) Multi-query attacks used a budget of 8 for reasonable computational cost across all models and attack settings for both sampling- and search-based attacks; (iii) The query scaling trends of specific LLMs and different attack settings were investigated with fixed budgets under 100 per input instance.
-  • Text embedding: Both PDWS and the DiscreteZOO attacks require sampling from a text embedding.
-The counter-fitted GloVe embedding (Mrkšić et al., 2016) was a default for DiscreteZOO, but GloVe is poor for representing specialized vocabulary in the biomedical domain (Wang et al., 2018).
-Therefore, here we chose two recent embedding models that offer an improved latent representation of text relations for the benchmarks: (i) CODER (Yuan et al., 2022), a BERT-structured embedding model contrastively pretrained to distinguish UMLS knowledge graph concepts; (ii) GTE-base (Li et al., 2023b) from constrastive learning of general web-scale text data.
-  • Hyperparameters: For single-query PDWS attacks, we tuned the primary hyperparameter $n$ within the interval of [-50, 50] using grid search with a step of 5 or 10 because the local maxima of ASR appear on the positive and negative sides.
-For multi-query PDWS attacks, hyperparameter $n$ was briefly tre-uned around its optimal value in the single-query attack.
+The following three aspects in the benchmarks are in our focus.
+  • Query budget: We used fixed query budgets ($B$) for three main types of attacks: (i) Single-query sampling-based attacks were used as the reference because the DiscreteZOO attack requires a minimum of 3 model queries; (ii) Multi-query attacks used a budget of 8 for reasonable computational cost across all models and attack settings for both sampling- and search-based attacks; (iii) The query scaling trends of specific LLMs and different attack settings were investigated with a series of query budgets under 100 per input instance.
+Here, we also included the deterministic versions of PDWS to sample the nearest and farthest elements with respect to an anchor point.
+The element was taken from the perturbation set and the distances (between entities) were calculated using cosine in a chosen text embedding.
+Their query-limited versions are refered to as $B$-nearest and $B$-farthest element sampling, respectively, where all $B$ nearest or farthest elements were used sequentially to attack models in query scaling studies.
+Detailed definitions are given in Appendix B.4.
+  • Text embedding: Both the PDWS and DiscreteZOO attacks require sampling from a text embedding.
+The counter-fitted GloVe embedding (Mrkšić et al., 2016) was the default for DiscreteZOO, but GloVe is poor for representing biomedical vocabulary (Wang et al., 2018).
+We chose two recent embedding models with improved latent representations of text relations for the benchmarks: (i) CODER (Yuan et al., 2022), a BERT-structured embedding model contrastively pretrained to distinguish UMLS knowledge graph concepts; (ii) GTE-base (Li et al., 2023b) from contrastive learning of general web-scale text data.
+  • Core hyperparameters: For single-query PDWS attacks, we tuned the hyperparameter $n$ within the interval of [-50, 50] using grid search with a step of 5 or 10 because the local maxima of ASR appear on the positive and negative sides.
+For multi-query PDWS attacks, hyperparameter $n$ was briefly re-tuned around its optimal value in the single-query attack.
 Most tuned hyperparameters fall within [-30, -5] and [5, 30].
@@
 5 Finding and assessing adversarial entities
-We conducted failure mode analysis using the attack statistics gathered across models and hyperparameter settings.
-The following analyses compare methods to identify adversarial entities through query-efficient attacks and uncover their characteristics, including the distribution in the embedding space and their effects on model explainability.
 Figure 2: Powerscaled DWS of adversarial distractors exhibits a two-regime effect at negative and positive $n$ values (see Eq.
-3) in model accuracy (top) and diversity index (bottom) of replacement entities in successful attacks.
-The observed similar behaviors are compared across models and datasets in (a) Flan-T5-xxl on MedQA-USMLE, (b) Palmyra-Med-20B on MedQA-USMLE, (c) Flan-T5-xl on MedMCQA, (d) Flan-UL2 on MedMCQA.
+2) in ASR (top) and diversity index (bottom) of replacement entities in successful attacks.
+Local maxima in ASR with a finite $n$ are also present in each regime.
+The vertical dashed line indicates the location of random sampling.
+The observed similar behaviors are compared across models and datasets in (a) Flan-T5-xxl on MedQA-USMLE, (b) Palmyra-Med-20B on MedQA-USMLE, (c) Flan-T5-xl on MedMCQA, (d) MedAlpaca-7B on MedMCQA.
 Disease and drug-mention questions are separated by colors.
-The vertical dashed line indicates the location of random sampling.
+The average prompt semantic similarity displayed in (e)-(h) is calculated for the successful attacks obtained from the corresponding attack settings in (a)-(d), respectively.
 {NiceTabular}
-@|l @ P1cm @ |P1.2cm @ P1.2cm @ P2.4cm @ |P1.2cm @ P1.2cm @ P2.4cm |@ MedQA-USMLE MedMCQA
+@|l @ P1.2cm @ |P1.2cm @ P1.2cm @ P2.6cm @ |P1.2cm @ P1.2cm @ P2.6cm |@ MedQA-USMLE MedMCQA
 Model (size) Entity
@@
 Diseases 0.396 10.1 15.7 / 13.1 0.408 15.7 25.5 / 17.7
-MedAlpaca-7B Drugs 0.474 7.4 9.3 / 10.3 0.433 11.6 13.2 / 12.7
-Diseases 0.488 9.9 13.3 / 10.7 0.436 13.5 18.1 / 14.2
+GPT-3.5 (175B) Drugs 0.438 9.1 11.2 / 10.7
+Diseases 0.453 8.6 11.5 / 8.8
+MedAlpaca-7B Drugs 0.479 7.2 8.2 / 9.6 0.448 13.9 16.2 / 15.9
+Diseases 0.498 8.4 12.5 / 9.2 0.434 14.8 19.1 / 13.7
+JSL-MedLM3-8Bv2 Drugs 0.827 10.4 10.2 / 11.8
+Diseases 0.771 9.5 18.5 / 11.3
 LM2-Medtuned-13B Drugs 0.371 16.4 18.1 / 17.8 0.568 15.3 17.3 / 18.1
@@
 For each model, the highest ASR for a type of entity-mention question in each dataset is bolded.
+The models with an asterisk () were only evaluated on one dataset to limit the computational cost.
 5.1 Two regimes of adversarial entities
@@
 The embedding space picture of the substitution attack in Fig.
-1a provides a geometric interpretation of powerscaled DWS to examine different types of adversarial examples by tuning $n$.
+1a provides a geometric interpretation of PDWS to examine different types of adversarial examples by tuning $n$.
 We use the Gini-Simpson index (Rao, 1982), $\eta_{\text{GS}}=1-\sum_{j}q_{j}^{2}$, to quantify the diversity of adversarial examples and a low value indicates less diversity.
 Here, $q_{j}\in(0,1]$ is the probability of occurrence of the entity $e_{j}$ among all instances of adversarial entities obtained from sampling, and $\sum_{j}q_{j}=1$.
-The diversity index indicates the reusability of adversarial entities which we use as a proxy indicator for the universal adversarial attack regime (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a).
-The dependence of the ASR and diversity on the hyperparameter $n$ in Fig.
-2 indicates two regimes: (i) the adversarial entities semantically close to ($n<0$) the key $k$ have a greater diversity; (ii) the adversarial entities semantically far from ($n>0$) the key $k$ are more reusable.
-TCES in these two regimes both can lead to an increase in successful attacks from random sampling (equivalent to $n=0$), as shown in Tables 5-5.1.
+A low diversity index indicates the reusability and is a proxy indicator for the universal adversarial attack regime (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a).
+Combining the ASR dependence on diversity and hyperparameter $n$ in Fig.
+2 reveals two regimes: (i) the adversarial entities semantically close to ($n<0$) the key $k$ have a greater diversity; (ii) the adversarial entities semantically far from ($n>0$) the key $k$ are more reusable.
+TCES in these two regimes both can lead to an increase in successful attacks from random sampling (equivalent to $n=0$), as shown in Tables 5-5.2.
 Fig.
@@
 Moreover, there exist local maxima of ASR in both regimes with finite $n$, which were discovered by hyperparameter tuning.
-We call these behaviors the two-regime effect and they were observed in almost all models evaluated here in varying magnitudes for both QA datasets, thanks to the controllable spatial coverage of sampling methods incorporating distance information.
+We call the behavior the two-regime effect and they were observed in almost all models evaluated here in varying magnitudes for both QA datasets, thanks to the controllable spatial coverage of sampling methods incorporating distance information.
 In both single-query and multi-query attacks, the benchmarks show that the $n<0$ regime of PDWS attacks tends to have a greater influence on model performance than the $n>0$ regime, although prominent counterexamples do exist (MedAlpaca-7B and Llama2-Medtuned-13B (for drug-mention questions), and Flan-T5-xl (for disease-mention questions) in Table 5).
@@
 Perturbations by powerscaled DWS (PDWS) in the $n<0$ and $n>0$ regimes are in blue and grey, respectively.
+Semantic distortion by adversarial entities
+Besides ASR, evaluating the semantic distortion is another way to quantify the effectiveness of adversarial attacks.
+We estimated the semantic distortion of entire prompts in successful attacks before and after entity substitution, corresponding to $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ in Definition 3.2.
+We calculated the average prompt semantic similarity (PSS) using SentenceTransformer (Reimers & Gurevych, 2019) with the RoBERTa-large model (all-roberta-large-v1).
+At each distinct $n$, the PSS $\in[-1,1]$ with 1 being the most similar, is averaged over all successful attack instances to obtain the average PSS, a quantitative measure of the semantic distortion.
+Its dependence on the hyperparameter $n$ for different models and datasets are shown in Fig.
+2.
 Adversarial entity characteristics
-The two-regime effect indicates that the success of the attack may come from two sources.
-At small distances, semantic proximity (Mozes et al., 2021) reduces the distinguishability between the key and the distractor, therefore increasing the chance of model failure.
-Whereas at larger distances, the attack succeeds potentially due to the obscurity of the distractor entities in the question context, which can also profoundly impact the model performance (Li et al., 2023a).
+The two-regime effect indicates that the success of the attacks may come from two sources that dominate at different locations on the same attack surface, which refers to entity substitution in the present work.
+  • At larger distances, the attack succeeds potentially due to the obscurity of the distractor entities in the question context, which can also profoundly impact the model performance (Li et al., 2023a).
 This explanation may be further supported by examining the diversity index, $\eta_{\text{GS}}$, which shows the dominance of a few replacement entities in successful attacks at large distances.
 We found these highly reusable adversarial entities by ranking their occurrences in successful attacks by powerscaled DWS.
-For drug names, the top-ranked adversarial entities include n-acetylglucosamine (an amino sugar and anti-inflammatory drug)777n-acetylglucosamine on RxList and technetium Tc 99m exametazime (a radiopharmaceutical and contrast agent)888technetium Tc 99m exametazime on RxList.
-For disease names, some of the chromosome deletion syndromes and rare diseases names are more reusable than others.
+For drug names, the top-ranked adversarial entities include n-acetylglucosamine (an amino sugar and anti-inflammatory drug)888n-acetylglucosamine on RxList and technetium Tc 99m exametazime (a radiopharmaceutical and contrast agent)999technetium Tc 99m exametazime on RxList.
+For disease names, some of the chromosome deletion syndromes and rare disease names are more reusable than others.
 These entities dominate the sampling results in the $n>0$ regime for all LLMs, indicating their transferability and universality.
+  • At small distances, semantic proximity (Mozes et al., 2021) reduces the distinguishability between the key and the distractor, therefore increasing the chance of model failure.
+Examples include similar drugs or diseases with insufficient details such as using diabetes as a distractor instead of the correct answer diabetes II (refers to type-II diabetes) can result in a successful attack.
+  • Adversarial distractors result in little semantic distortion at the prompt level (see Fig.
+2e-h).
+The experiemnts on the MedMCQA dataset yields somewhat lower average PSS than MedQA-USMLE because of the longer format of the latter dataset, leading to sightly larger distortions by entity substitution.
+Adversarial entities obtained in the $n<0$ regime results in slightly higher average PSS compared to those in the $n>0$ regime, while there is also a two-regime effect – the average PSS tends to be the lowest around $n=0$ and higher when $n$ departs further from 0.
+The trend is consistent for different models and datasets demonstrated in Fig.
+2.
+5.2 Scaling characteristics from adversarial entities
+Size is not all for scaling model robustness.
+We compare model robustness using single-query ASR as the metric in Table 5 and Fig.
+3.
+For the Flan-T5 series of GLMs (Chung et al., 2024) and Flan-UL2 (Tay et al., 2022), the single-query ASR is more pronounced in smaller and less performant models, indicating the improvement of robustness as the model’s size grows.
+However, an opposite trend of robustness is seen in BLMs.
+It is worth noting that although Palmyra-Med-20B (Kamble & Alshikh, 2023) has about three times the number of parameters than MedAlpaca-7B (Han et al., 2023), the larger model is noticeably more sensitive to entity perturbation.
+In terms of ASR, the BLMs are comparable to the GLMs in each type of questions evaluated on.
+Another observation is that the ASR for models is asymmetrical under PDWS attacks.
+For most of the models, sampling adversarial entities with inverse-distance-scaled weights ($n<0$ in Eq.
+2) induces more performance drop than sampling with distance-scaled weights ($n>0$ in Eq.
+2), yet prominent counterexamples exist.
+The attack performance is further investigated in the sequential, multi-query setting (budget = 8) in Table 5.2, where we also compared sampling- against search-based DiscreteZOO attacks.
+Here, PDWS attacks in the $n<0$ regime are still largely the most effective.
 {NiceTabular}
-@|l @ P1cm @ |P1cm @ P1.8cm @ P1.3cm @ P1.3cm @ |P1cm @ P1.8cm @ P1.3cm @ P1.3cm |@ MedQA-USMLE (ASR %) MedMCQA (ASR %)
+@|l @ P1.1cm @ |P0.9cm @ P2cm @ P1.3cm @ P1.5cm @ |P0.9cm @ P2cm @ P1.3cm @ P1.5cm |@ MedQA-USMLE (ASR %) MedMCQA (ASR %)
 Model (size) Entity
@@
 Diseases 25.8 34.9 / 26.8 25.7 24.0 39.2 49.5 / 31.6 41.3 37.0
-MedAlpaca-7B Drugs 20.3 22.8 / 24.5 22.8 20.5 24.9 34.2 / 27.5 42.6 36.8
-Diseases 25.2 30.1 / 21.6 26.5 25.2 31.2 41.7 / 25.0 40.5 38.1
+MedAlpaca-7B Drugs 16.8 21.1 / 22.3 22.8 20.5 27.9 37.1 / 30.1 42.6 36.8
+Diseases 21.7 30.3 / 24.0 26.5 25.2 29.8 43.9 / 30.8 40.5 38.1
 Palmyra-Med-20B Drugs 30.1 31.4 / 31.3 29.1 29.6 28.5 37.3 / 31.9 35.4 35.0
@@
 Table 2: Multi-query (here budget = 8) attack success rates (ASR %) on various LLMs by perturbing the drug and disease entity mentions in MedQA-USMLE and MedMCQA datasets.
-The PDWS attack uses the embedding from CODER and DiscreteZOO (DZOO) attacks use either CODER or GTE-base to calculate the cosine distance.
+The PDWS attack uses the embedding from CODER and DiscreteZOO (DZOO) attacks use either CODER or GTE-base embedding.
 For each model, the highest ASR for a type of entity-mention question in each dataset is bolded.
-5.2 Scaling characteristics from adversarial entities
-Size is not all for scaling model robustness.
-We compare model robustness using single-query ASR as the metric in Table 5 and Fig.
-3.
-For the Flan-T5 series of GLMs (Chung et al., 2024) and Flan-UL2 (Tay et al., 2022), the single-query ASR is more pronounced in smaller and less performant models, indicating the improvement of robustness as the model’s size grows.
-However, an opposite trend of robustness is seem in BLMs.
-It is worth noting that although Palmyra-Med-20B (Kamble & Alshikh, 2023) has about three times the number of parameters than MedAlpaca-7B (Han et al., 2023), the larger model is much more sensitive to entity perturbation.
-In terms of ASR, the BLMs are comparable to the GLMs in each type of questions evaluated on.
-Another observation is that the ASR for models is asymmetrical under PDWS attacks.
-For most of the models, sampling adversarial entities with inverse-distance-proportional weights ($n<0$ in Eq.
-3) induces more performance drop than sampling with distance-proportional weights ($n>0$ in Eq.
-3), yet prominent counterexamples exist.
-The attack performance is further investigated in the sequential, multi-query setting (budget = 8) in Table 5.1, where we also compared sampling- against search-based DiscreteZOO attacks.
-Here, PDWS attacks in the $n<0$ regime are still largely the most effective.
-Figure 4: Example scaling curves of the query budget against ASR for disease-mention questions with (a) Flan-T5-xl on MedMCQA dataset and (b) MedAlpaca model on MedQA-USMLE dataset.
+Figure 4: Example scaling curves of the query budget against ASR for disease-mention questions with (a) Flan-T5-xl model on MedMCQA dataset and (b) MedAlpaca model on MedQA-USMLE dataset.
 The curves are generated using TCES of disease names with the methods described in the legends.
@@
 Executing the random sampling (RS) attack doesn’t require a text embedding, while the other attacks were evaluated with the CODER or GTE-base as text embedding.
-Sampling in specialized text embedding has favorable query scaling over blackbox gradient attacks.
-A query scaling study of adversarial attacks investigates attack success with varying query budgets per input instance (Shukla et al., 2021).
-It is a comprehensive evaluation of attack methods, but not yet frequently adopted in most publications, partly due to the computational cost.
-Fig.
-4 compares the query scaling of both sampling- and search-based methods for TCES.
-For low-query-budget attacks (Fig.
-1a), each scaling curve exhibits a rapid rise phase and a plateau phase, separated by an inflection point.
-The key message here is that the advantage of PDWS attacks over DiscreteZOO exists with a choice of text embedding.
-It is greater in specialized text embeddings but is reduced or diminished in general ones thanks to the difference in neighborhood structures.
-We expand the analysis in the following in three directions.
-  • Query scaling depends on text embedding: Changing the text embedding affects the query scaling behavior of both PDWS and DiscreteZOO attacks.
-Fig.
-4 shows that the influence of the embedding choice on PDWS is more pronounced in the rapid rise phase, while it influences the DiscreteZOO primarily in the plateau phase.
-Specifically, changing the text embedding from domain-specialized CODER (Yuan et al., 2022) to the general GTE-base (Li et al., 2023b), the performance advantage of PDWS over DiscreteZOO attacks shrunk in Flan-T5-xl (Fig.
-4a) and inverted in MedAlpaca-7B (Fig.
-4b).
-  • Sampling hyperparameter controls scaling curve shape in PDWS attacks: The hyperparameter $n$ changes the inflection point of the scaling curve.
-When the query budget is past the inflection point of the curve, random sampling starts to surpass PDWS ($n>0$) in ASR.
-This is a consequence of the distribution of adversarial entities in the embedding space and the regional bias derived from the sampling weights.
-The examples in Fig.
-4 show that the largest performance gap between the PDWS and DiscreteZOO attacks appears at around the inflection point of the respective query scaling curve.
-  • Gradient quality affects query scaling: The query complexity in blackbox gradient estimation is a long-standing theoretical topic in derivative-free optimization (L’Ecuyer & Yin, 1998) and online learning (Duchi et al., 2015).
-Nevertheless, in application settings such as adversarial machine learning, how gradient quality affects query scaling remains less studied for text data.
-Gradient estimation consumes a part of the total query budget, leading to a tradeoff between the gradient quality and terminal ASR attainable in the query scaling curve.
-We empirically explored the query scaling behavior of the DiscreteZOO attack in Appendix E with different candidate points in the multi-point gradient estimation, which in effect alters the gradient quality.
-The results show that having too few candidate points (i.e.
-small $M$ in Eq.
-6) leads to an early onset of the plateau phase, thereby limiting the highest attainable ASR within the total query budget, whereas having many candidate points can lead to a slower growth of ASR in the rapid rise phase of the query scaling curve.
+In (c) and (d), query scaling of the attacks based on $B$-nearest and $B$-farthest element sampling ($B$ is query budget) for the same datasets and models as in (a) and (b) are compared with PDWS.
 A 73-year-old man presents to the outpatient clinic complaining of chest pain with exertion.
@@
 Figure 5: (Top) An entity substitution attack using n-acetylglucosamine as the replacement entity for metoprolol creates an adversarial distractor.
-(Bottom) Heatmaps of token-wise Shapley values for a question before and after the adversarial attack on option C.
+(Bottom) Heatmaps of token-wise Shapley values for a question before and after the adversarial attack on choice C.
 The model prediction changes from the correct choice of B (unperturbed) to the incorrect choice of D (adversarially perturbed).
+Sampling attacks has improved query scaling over blackbox gradient attacks in specialized text embedding.
+A query scaling study of adversarial attacks investigates attack success with varying query budgets per input instance (Shukla et al., 2021).
+It is a comprehensive evaluation of attack methods.
+Fig.
+4 compares the query scaling of both sampling- and search-based methods for TCES.
+For low-query-budget attacks (Fig.
+1a), each scaling curve exhibits a rapid rise phase and a plateau phase, separated by an inflection point.
+The key message here is that the advantage of PDWS attacks over DiscreteZOO depends on the model and the choice of text embedding.
+The PDWS attacks are overarchingly more effective for GLMs than DiscreteZOO, while for BLMs, the advantage is reduced.
+Moreover, using a specialized text embedding further boosts the advantage of PDWS attacks compared with a general text embedding due to the difference in embedding space neighborhood.
+We expand the analysis in the following in three directions.
+  • Query scaling depends on text embedding: Changing the text embedding affects the query scaling behavior of both PDWS and DiscreteZOO attacks.
+Fig.
+4 shows that the influence of the embedding choice on PDWS is more pronounced in the rapid rise phase, while it influences the DiscreteZOO primarily in the plateau phase.
+Specifically, changing the text embedding from domain-specialized CODER (Yuan et al., 2022) to the general GTE-base (Li et al., 2023b), the performance advantage of PDWS over DiscreteZOO attacks shrunk in Flan-T5-xl (Fig.
+4a) and inverted in MedAlpaca-7B (Fig.
+4b).
+  • Sampling hyperparameter controls scaling curve shape in PDWS attacks: The hyperparameter $n$ changes the inflection point of the scaling curve.
+When the query budget passes the inflection point of the curve, random sampling starts to surpass PDWS ($n>0$) in ASR.
+This is a consequence of the distribution of adversarial entities in the embedding space and the regional bias derived from the sampling weights.
+The examples in Fig.
+4 show that the largest performance gap between the PDWS and DiscreteZOO attacks appears at around the inflection point of the respective query scaling curve.
+  • Gradient quality affects query scaling in DiscreteZOO attacks: The query complexity in blackbox gradient estimation is a long-standing theoretical topic in derivative-free optimization (L’Ecuyer & Yin, 1998) and online learning (Duchi et al., 2015).
+Nevertheless, how gradient quality affects query scaling remains less studied for adversarial attacks employing gradient-based discrete optimization.
+Gradient estimation consumes a large part of the total query budget, leading to a tradeoff between the gradient quality and terminal ASR attainable (at max budget) in the query scaling curve.
+We empirically explored the query scaling behavior of the DiscreteZOO attack in Appendix E with different candidate points in the multi-point gradient estimation, which in effect alters the gradient quality.
+The results show that having too few candidate points (i.e.
+small $M$ in Eq.
+5) leads to an early onset of the plateau phase, thereby limiting the highest attainable ASR, whereas having many candidate points can lead to a slower growth of ASR in the rapid rise phase of the query scaling curve.
+  • Deterministic sampling attacks are competitive proxies of their probabilistic counterparts.
+Performance comparison between $B$-nearest and $B$-farthest element sampling, the deterministic and query-limited versions of PDWS (corresponding to the $n<0$ and $n>0$ regimes) are given in Fig.
+4c-d for the two example cases studied for the query scaling.
+The results show that the deterministic sampling attacks are already competitive against DiscreteZOO, and also close in performance to PDWS in both specialized and general text embeddings.
+The difference is that the two types of deterministic sampling attacks examine only predefined regions in the embedding space and cannot explore the entire attack surface as PDWS would through tuning the $n$ hyperparameter.
 5.3 Adversarial entities manipulate explainability
@@
 6 Discussion and future work
-The present study illustrates that distracting information introduced through QA distractors can lead to noticeable performance degradation in billion-parameter LLMs.
-TCES with only a single entity modification step combined with adversarial sampling methods and appropriate text embeddings is a feasible way for collecting adversarial entity data.
-Moreover, TCES ensures that the replacement entity appears natural to humans and is less likely to be detected using grammar check or topical filtering (Morris et al., 2020a).
-The setting simulates realistic scenarios where the “attacks" may be initiated by unsuspecting users, such as biomedical professionals using an LLM-based clinical decision support system (Liu et al., 2023).
+The present study illustrates that TCES is a compact yet effective way to construct adversarial distractors, which leads to significant performance degradation in LLMs.
+Our results suggest that to minimize model queries, using a budget of up to the inflection point on the query scaling curve is the most cost-effective.
+The perturbed texts produced by TCES appear natural and are less likely to be detected using grammar check or topical filtering (Morris et al., 2020a).
+Therefore, the setting can simulate realistic scenarios where the “attacks” may be initiated by unsuspecting users, such as healthcare professionals using an LLM-based clinical decision support system (Liu et al., 2023).
 In the examples illustrated in Fig.
 5 and Appendix E, adversarial distractors can lead to disease misdiagnosis or misprescription of medication, which are detrimental scenarios that may occur in algorithmic decision-making in biomedicine using LLMs.
-Further improvement on the attack performance may be achieved using learned samplers through rejection mechanisms (Narasimhan et al., 2024) or by finetuning the text embedding to improve the query efficiency of sampling.
-Our approach may be integrated into interactive platforms for adversarial testing in open-ended human-machine conversation (Chao et al., 2023) for online model auditing.
+Further improvement on the sampling attack performance may be achieved using learned samplers through rejection mechanisms (Narasimhan et al., 2024) or by finetuning the text embedding to improve the query efficiency of sampling.
+Alternatively, the semantic distance used for PDWS may be replaced with a concept distance (Choi et al., 2016) to improve attack performance on general text embeddings.
+Our approach may be integrated into interactive platforms for adversarial data collection or online monitoring systems for open-ended human-machine conversation (Chao et al., 2023).
 Future research may focus on how to make LLMs more reliable through prompt engineering (Si et al., 2023; Nori et al., 2023) or leveraging retrieval from nonparametric knowledge bases (Weikum et al., 2021; Soman et al., 2023) or text sources (Wang et al., 2023b) to improve model generalization to long-tail entities.
-Besides vulnerability identification, the current work motivates adversarial defense strategies based on generalized distances to establish a tiered system for validating adversarial robustness in user-facing and domain-oriented applications.
-Broader Impact Statement
-The scenarios considered in this work using public datasets relate to the potential real-world deployment of LLMs in biomedical education and clinical decision support.
-Our results show that the lack of robustness assessment or adversarial testing methods in high-stakes domains may incur serious consequences because of naturally looking adversarial examples.
-The design and deployment of adversarial testing should be specific to and carried out at every stage of the model development cycle, paired with tools such as adversarially robust model explainability methods for understanding model behavior.
+Besides vulnerability identification, the current work motivates adversarial defense strategies based on generalized distances (La Malfa et al., 2020) to establish a tiered system for robustness assessment in user-facing and domain-oriented applications to mitigate catastrophic failures.
+Acknowledgments
+R.P.X.
+thanks N.
+Rethmeier for helpful technical discussions.
+A.J.L.
+acknowledges funding from NIH T32GM067547, the UCSF Discovery Fellows Program, and the Achievement Rewards for College Scientists (ARCS) Fellowship.
+R.A.-A.
+would like to acknowledge funding from Weill Neurohub.
 Appendix A Adversarial distractor generation (General case)
@@
 For the entities in the perturbation set -$\mathcal{E}_{\textrm{-}k}$, this indicates that only
-$e_{j}=\mathrm{argmax}\,\mathcal{H}_{\textrm{-}k,\uparrow}=\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmax}}\,h(k,e_{j})$ (13)
+$e_{j}=\mathrm{argmax}\,\mathcal{H}_{\textrm{-}k,\uparrow}\equiv\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmax}}\,h(k,e_{j})$ (12)
 is selected with certainty at the deterministic limit.
+The equivalence ($\equiv$) is used because the two sets -$\mathcal{H}_{\textrm{-}k,\uparrow}$ and $h(k,e_{j})$ $(j=1,...,m)$ have the same elements.
+Eq.
+12 also defines the farthest element sampling.
 B.2 The limit of $n\rightarrow-\infty$
@@
 As noted in the footnote of the main text, this limit is not attained since $k\notin\mathcal{E}_{\textrm{-}k}$.
+Nevertheless, we can define the deterministic sampling of the nearest entity to $k$ in a similar vein as Eq.
+12,
+$e_{j}=\mathrm{argmin}\,\mathcal{H}_{\textrm{-}k,\uparrow}\equiv\underset{e_{j}\in\mathcal{E}_{\textrm{-}k}}{\mathrm{argmin}}\,h(k,e_{j}),$ (13)
+which we refer to as nearest element sampling.
 B.3 The case of $n=0$
 When $n=0$, it is straightforward to show that $p_{j}|_{n=0}=1/m$, regardless of $j$ using the definition in Eq. 2. This scenario therefore corresponds to uniform random sampling.
+B.4 Query-limited deterministic sampling
+Given a query budget $B$, construct the ordered sets -$\mathcal{H}_{\textrm{-}k,\uparrow}^{j,m}=\{h_{j,\uparrow},...,h_{m,\uparrow}\}$ ($j=1,...,B$), and assuming $B\leq m$, we know that -$\mathcal{H}_{\textrm{-}k,\uparrow}^{j,m}\subseteq\mathcal{H}_{\textrm{-}k,\uparrow}\equiv\mathcal{H}_{\textrm{-}k,\uparrow}^{1,m}$ from the definition of -$\mathcal{H}_{\textrm{-}k,\uparrow}$ in Appendix B.1.
+We use $B$-nearest element sampling to refer to the deterministic sampling carried out sequentially where at the $j$th ($j=1,...,B$) query, the sampled entity $e_{j}=\mathrm{argmin}\,\mathcal{H}_{\textrm{-}k,\uparrow}^{j,m}$.
+Similarly, we can construct the ordered sets -$\mathcal{H}_{\textrm{-}k,\uparrow}^{1,m-j+1}=\{h_{1,\uparrow},...,h_{m-j+1,\uparrow}\}$ ($j=1,...,B$).
+We use $B$-farthest element sampling to refer to the sequential deterministic sampling where at the $j$th ($j=1,...,B$) query, the sampled entity $e_{j}=\mathrm{argmax}\,\mathcal{H}_{\textrm{-}k,\uparrow}^{1,m-j+1}$.
+When $B=1$, the query-limited versions reduces to the farthest and nearest element sampling described earlier in Appendix B with only a single query.
 Prompt format Answer the following question without explanation.
@@
 According to Eq.
-6, for each entity to attack, we randomly sampled two, five, and ten candidate points in its neighborhood in the embedding space for gradient estimation.
+5, for each entity to attack, we randomly sampled two, five, and ten candidate points in its neighborhood in the embedding space for gradient estimation.
 The attack on each input instance proceeds with a fixed total of LLM queries (total query budget) as indicated on the horizontal axis.
@@
 5 in the main text.
-For each example, the text box on top (the prompt formatting removed) shows the original options on the left side and the perturbed options on the right side.
+For each example, the text box on top (the prompt formatting removed) shows the original choices on the left side and the perturbed choices on the right side.
 The entity of interest before and after perturbation is colored in red and blue, respectively.